### v1

# Useful entanglement can be extracted from noisy graph states

Abstract
Cluster states and graph states in general offer a useful model of the stabilizer formalism and a path toward the development of measurement-based quantum computation. Their defining structure – the stabilizer group – encodes all possible correlations which can be observed during measurement. Those outcomes which are compatible with the stabilizer structure make error correction possible. Here, we leverage both properties to design feasible families of states that can be used as robust building blocks of quantum computation. This procedure reduces the effect of experimentally relevant noise models on the extraction of smaller entangled states from the larger noisy graph state. In particular, we study the extraction of Bell pairs from linearly extended graph states – this has the immediate consequence for state teleportation across the graph. We show that robust entanglement can be extracted by proper design of the linear graph with only a minimal overhead of the physical qubits. This scenario is relevant to systems in which the entanglement can be created between neighboring sites. The results shown in this work may provide a mathematical framework for noise reduction in measurement-based quantum computation. With proper connectivity structures, the effect of noise can be minimized for a large class of realistic noise processes.
1 Introduction
The discovery of cluster states – states of qubits with grid-like entanglement structure – provided a new perspective for quantum computation, better applicable for some systems [1, 2]. The original design consisted of qubits arranged in a two-dimensional square lattice structure, where the neighboring qubits would be entangled by means of a controlled-$Z$ gate (or a specific quantum system equivalent thereof). Unlike the circuit-based understanding of quantum computation, the input state is constant, and the only available class of operations is the sequential measurement of individual qubits. It is the act of measurement that feeds the information into the system and performs all of the needed transformations [3]. Generalizations into different connectivity structures – the (hyper)graph states – soon followed, sharing the same core properties with the original cluster states [4, 5, 6, 7] and providing a framework for further research into the entanglement properties of multiqubit states [8, 9].
The most basic operation in this context is no processing at all: a simple transfer or teleportation [10, 11] of one qubit state into another place. It is a basic building block of measurement-based quantum computation, on top of which more involved operations are built. In the basic scenario, this can be performed by sequential measurement of Pauli-$X$ operators on a path between the initial and end qubits (the terminal qubits in this article nomenclature) and Pauli-$Z$ on the qubits neighboring the path. Intuitively, the $Z$-basis measurement excises the entanglement contained in the path from the rest of the cluster state and the $X$-basis measurement along the path makes the terminal qubits correlated: final measurement of one of them fuses the two together, and the surviving qubit shares correlations with a distant part of the graph [3]. This process can also be understood as a Bell pair generation along the path: local operations do commute, and a measurement along the path on its own does produce Bell-like correlations across the terminal qubits, which can be used directly for information transfer. All of this works, provided that no noise was present during the state preparation and measurement – but this is rarely the case. Thus, a noise-detecting or -correcting process is needed, and research in this direction has been done [12, 13, 14, 15, 16].
In general, entanglement is required for quantum computation, and preparation of a Bell pair or GHZ state is an operational primitive of many quantum algorithms. Extracting those states from larger graph states was discussed in [17, 18, 14, 19, 20].
In this article we analyze experimentally motivated noise models and determine what can be done (to the initial noisy graph state structure and the measurement pattern) in order to mitigate the noise effects. This approach does not translate into full-fledged error correcting codes, but still uses the properties of the stabilizer associated with the graph state; in the ideal case, some outcomes are perfectly (anti)correlated, and any deviation indicates the presence of noise. As mentioned in the previous paragraph, teleportation is equivalent to a Bell pair preparation; in order to simplify the calculations, we analyze the entanglement quality of the prepared state after the internal qubits have been measured.
The article is structured as follows. In Section 2 we briefly introduce graph states and the stabilizer formalism. In Section 3 we explain the measurement scheme which we use to generate Bell pairs between certain nodes or for teleportation. We introduce noise models in Section 4 and analyze how to deal with their effects in Section 5. In Section 6 we show how to extend the presented methods to generate GHZ states. We conclude our work in Section 7.
2 Graph state preliminaries
Figure 1: a) The square graph $G$. The associated state |$\ket{G}$, as any graph state, possesses an associated structure of the stabilizer $S_{G}$ composed of specific strings of Pauli operators. Expectation values of any Pauli string $O$ is nonzero if and only if ±$\pm O\in S_{G}$. $X_{1}X_{3}$ is a stabilizer operator of |$\ket{G}$ and therefore it’s expectation value is 1.
b) Graphs corresponding to the state defined in a) being altered by noise – here, the noise may remove certain edges (dashed lines). Noise processes change the stabilize group (e.g. by graph modification shown here), which allows for verifiable preparation.
In this section, we present a short introduction to the class of graph states and the stabilizer formalism. Readers familiar with the topic may skip directly to the next section.
A graph $G=(V,E)$ is defined by a set $V$ of vertices and a set of edges $E\subset V^{2}$: each $\{i,j\}=:e\in E$ denotes a connection between two vertices $i,j\in V$. Graph states are multi-qubit quantum states, where the vertices and edges of the graph $G=(V,E)$ represent qubits and entangling gates, respectively. The state |$\ket{G}$, corresponding to a graph $G$ is defined as
|$\ket{G}=\mathchoice{}{}{}{}\prod_{\{i,j\}\in E}\mathchoice{}{}{}{}\operatorname{CZ}_{i,j}\ket{+}^{\otimes\lvert V\rvert},$ (1)
where $\operatorname{CZ}$ denotes a standard controlled-$Z$ unitary operation
[…]
Manifestly, all of controlled-$Z$ gates are diagonal in the computation basis, and hence commute – the product in the above equation does not depend on the order of operations.
Consider a four-vertex graph shown in Figure 1a: it has a vertex set $V=\{1,2,3,4\}$ and an edge set $E=\{\{1,2\},\{2,3\},\{3,4\},\{4,1\}\}$. The corresponding graph state is given by
|$\ket{G}=\operatorname{CZ}_{1,2}\operatorname{CZ}_{2,3}\operatorname{CZ}_{3,4}\operatorname{CZ}_{4,1}\ket{+}^{\otimes 4}.$ (3)
The graph states can also be described using the so-called stabilizer operators, providing a description of the state in question in terms of its correlation structure. Consider an initial state |$\ket{+}^{\otimes|V|}$: since it is a pure product state, there are no correlations between different sites. However, it is an eigenstate for each $X_{i}$ for $i\in V$. Let us denote the product of controlled-$Z$ operators in Equation 1 by $U$; then
|$\ket{G}$ $=U\ket{+}^{\otimes\lvert V\rvert}=UX_{i}\ket{+}^{\otimes\lvert V\rvert}$ (4)
$=\underbrace{UX_{i}U^{\dagger}}_{g_{i}}\underbrace{U\ket{+}^{\otimes\lvert V\rvert}}_{\ket{G}}=g_{i}\ket{G}.$ (5)
The operator $g_{i}$ can be shown to have the form of
$g_{i}=X_{i}\mathchoice{}{}{}{}\prod_{j\in\operatorname{\mathcal{N}}(i)}\mathchoice{}{}{}{}Z_{j},$ (6)
where $X_{i}$, $Y_{i}$, and $Z_{i}$ denote the Pauli matrices acting on the $i$-th qubit and $\operatorname{\mathcal{N}}(i)$ is the neighbourhood of $i$. As shown above, any graph state |$\ket{G}$ defined by Equation 1 is an eigenstate (to the eigenvalue of $+1$) of all $g_{i}$:
$g_{i}\ket{G}=\ket{G}$ (7)
While the $(+1)$-eigenspace of each of the operators $g_{i}$ is highly degenerated, taken together they fully define the state |$\ket{G}$: it is the $(+1)$-eigenstate of all $g_{i}$. Operators defined by Equation 7 commute; hence, they generate an Abelian subgroup of the group of Pauli operators, called the stabilizer $S$ of |$\ket{G}$:
$S=\left\{\prod_{i\in I}g_{i}:I\subset V\right\}.$ (8)
The elements of the group $S$, due to the compact form of the generators (Equation 6), can be described with the help of low-dimensional matrices of integers modulo $2$ [21]. This representation is helpful in direct calculations involving the stabilizer.
The exemplary graph state shown in Figure 1a has stabilizer operators $g_{1}=X_{1}Z_{2}\operatorname{\mathds{1}}_{3}Z_{4}$, $g_{2}=Z_{1}X_{2}Z_{3}\operatorname{\mathds{1}}_{4}$, $g_{3}=\operatorname{\mathds{1}}_{1}Z_{2}X_{3}Z_{4}$, and $g_{4}=Z_{1}\operatorname{\mathds{1}}_{2}Z_{3}X_{4}$, as well as all products composed from these operators.
The elements of $S$ form a complete description of all possible correlations of local measurements of graph states: if a Pauli string [^1] Here, $o_{i}$ denotes one of the Pauli operators $\{X_{i},Y_{i},Z_{i}\}$ acting on the $i$-th qubit. $O=o_{i_{1}}o_{i_{2}}\ldots$ (or its negative, $-O$) appears in $S$, its expectation value on |$\ket{G}$ is equal to $+1$ (respectively, $-1$), and if ±$\pm O\notin S$, $\langle O\rangle_{G}=0$.
For example, for the graph state shown in Figure 1a, the string $O=X_{1}X_{3}=g_{1}g_{3}$ is a stabilizer operator of |$\ket{G}$. Using Equation 7 it can be seen that $\langle O\rangle_{G}=1$. The string $O^{\prime}=X_{1}Z_{3}\notin S_{G}$, hence $\langle O^{\prime}\rangle_{G}=0$. In this paper we are mostly using Pauli strings which contain only Pauli $X$ operators [21]. However, the results can be generalized to other Pauli strings.
The weighted graph states [22] are a related class of quantum states, for which controlled-phase operators are used instead of controlled-$Z$, with the phases possibly different for each edge of the graph $G$. If the phase associated with the edge $\{i,j\}$ is $\varphi_{i,j}$, the resulting state is defined by a modified version of Equation 1,
|${\ket{G_{\varphi}}=\mathchoice{}{}{}{}\prod_{\{i,j\}\in E}\mathchoice{}{}{}{}\operatorname{C\Phi}_{i,j}(\varphi_{i,j})\ket{+}^{\otimes\lvert V\rvert},}$ (9)
where $\operatorname{C\Phi}(\varphi)=\operatorname{diag}(1,1,1,\exp(i\varphi))$ on the $\{i,j\}$ subsystem.S uch states arise naturally in systems with Ising-like interactions, and offer an useful representation of noisy graph state preparation using such a scheme. Unfortunately, the analogue of procedure described by Equation 5 does not produce a Pauli string, and hence a simple stabilizer description of correlations generally does not exist.
For a standard graph state |$\ket{G}$ all phases $\varphi_{i,j}$ above are equal to $\pi$. This class of states naturally arises in systems with an Ising-type interaction pattern, and some physical realizations of graph states prepare the unweighted graph states in this exact way.
Many known families of states with structured entanglement can be represented as graph states. For instance, the $N$-qubit Greenberger–Horne–Zeilinger (GHZ) state [23], defined as
|$\ket{GHZ}=\frac{1}{\sqrt{2}}(\ket{0_{1}0_{2}\dots 0_{N}}+\ket{1_{1}1_{2}\dots 1_{N}})$ (10)
is local unitary equivalent to the graph state of the fully connected $N$ vertex graph (or, equivalently, by the $N$ vertex star graph). In measurement-based quantum computing, cluster states, corresponding to grid graphs, are relevant [3], and intermediate-size cluster states have recently been realized on various platforms [24, 25]. A detailed discussion of graph states, including weighted graph states, can be found in [5, 26].
3 Correlations and measurement
It is possible to extract an entangled pair of qubits $(i,j)$ from a graph state |$\ket{G}$ by means of local Pauli measurements, if there exists a path between $i$ and $j$ through the graph $G$ [20, 3]. In general, there are multiple ways to perform this task – and if noise is present, the entanglement quality of the resulting two-qubit state depends on the choice.
Consider a graph state |$\ket{G}$ that undergoes a local sequential measurement process involving Pauli operators. The state of the unmeasured qubits is completely characterized by the measurement pattern along with the outcomes; the remaining correlations stem from the stabilizer operators of |$\ket{G}$ consistent with the measurement pattern.
To see the structure of these correlations, let us choose a measurement pattern on a subset of qubits $I$: for each qubit $i\in I$, a local Pauli $o_{i}\in\{X_{i},Y_{i},Z_{i}\}$ is chosen. Then, each qubit $i$ is measured sequentially in the eigenbasis of the associated operator $o_{i}$. As mentioned previously, the expectation value of a Pauli string $O=\prod_{i\in I}o_{i}$ can be determined: if $O\in S_{G}$, it implies $\expectationvalue{O}_{G}=1$. The case of $-O\in S_{G}$ differs only by the sign, and if neither ±$\pm O\in S_{G}$, the expectation value is 0.
Let us concentrate on the positive sign case (the negative sign does not change much in reasoning) and consider an actual preparation of |$\ket{G}$ and subsequent measurement according to $\{o_{i}\}_{i\in V}$: the measurement of $o_{i}$ on the $i$-th qubit has an outcome of $s_{i}$. If $\langle O\rangle_{G}=1$, the outcomes of the sequential measurement must reflect that: the signs of the measurement results must multiply to $+1$. Any other outcome pattern would decrease the magnitude of the experimental expectation value and is therefore incompatible with the graph state correlations.
This reasoning applies as well if the sequential measurement is stopped at any point and resumed afterward: if $O=\prod_{i\in I}o_{i}$ is a stabilizer, and only the qubits $I^{\prime}\subset I$ were measured, the sign structure of the further measurement of $I\setminus I^{\prime}$ can be predicted. Thus, by the action of partial measurement, correlation in the remaining qubits is induced.
The following Lemma captures this line of thought: the end correlations are defined by stabilizer operators consistent with the measurement scheme. Note that the measurement scheme does not itself have to be a stabilizer operator: only part of it has to extend to one, and this is always possible (see Lemma 3 in Appendix A).
Lemma 1.
Let |$\ket{G}$ be a graph state determined by the graph $G=(V,E)$; we denote the stabilizer of |$\ket{G}$ by $S$. If a subset $I$ of qubits is measured in such a way that a qubit $i\in I$ is measured in the eigenbasis of $o_{i}\in\{X_{i},Y_{i},Z_{i}\}$, we encode this measurement scheme as the Pauli string $O=\prod_{i\in I}o_{i}$. The qubits of $I$ are measured independently and sequentially, with the measurement outcome of $o_{i}$ denoted by $s_{i}=\pm 1$.
Let us now write a stabilizer operator $Q\in S$ of |$\ket{G}$ as a Pauli string: $Q=\pm\prod_{j\in J}q_{j}$. The post-measurement state |$\ket{\psi}$ is determined by those elements of the stabilizer, which are consistent with the measurement pattern $O$ such that $Q|_{I\cap J}=O|_{I\cap J}$. The stabilizer operators of the post-measurement state |$\ket{\psi}$ can be computed from such operators $Q$ as
$Q^{\prime}:=\pm\prod_{i\in I\cap J}s_{i}\prod_{j\in J\setminus I}q_{j}.$ (11)
where the global sign is equal to the global sign of $Q$.
Proof.
Consider the stabilizer operator $Q$ defined as stated in the Lemma. Then assume that the measurement outcomes of each of the qubits $i\in I\cap J$ are $s_{i}$. Then, the projective measurement operators of these qubits can be written as $\frac{1}{2}\left(\mathbbm{1}_{i}+s_{i}q_{i}\right)$. [^2] For instance, the projection operator associated with the $X+$ outcome is |$\ket{X+}\bra{X+}=\frac{1}{2}\left(\mathbbm{1}+X\right)$. Since the projection operators commute with $q_{j}$ themselves, the following chain of equalities holds:
|$\ket{\psi}$ $=C\prod_{i\in I\cap J}\frac{\mathbbm{1}_{i}+s_{i}q_{i}}{2}\underbrace{\ket{G}}_{\leavevmode\nobreak\ \leavevmode\nobreak\ =Q\ket{G}}$ (12)
$=C\prod_{i\in I\cap J}\frac{\mathbbm{1}_{i}+s_{i}q_{i}}{2}\left(\pm\prod_{j\in J}q_{j}\ket{G}\right)$
=±$=\pm\prod_{j\in J\setminus I}q_{j}\prod_{i\in I\cap J}s_{i}\underbrace{\left(C\prod_{i\in I\cap J}\frac{\mathbbm{1}_{i}+s_{i}q_{i}}{2}\ket{G}\right)}_{\leavevmode\nobreak\ \leavevmode\nobreak\ =\ket{\psi}}$
=(±$=\underbrace{\left(\pm\prod_{i\in I\cap J}s_{i}\prod_{j\in J\setminus I}q_{j}\right)}_{\leavevmode\nobreak\ \leavevmode\nobreak\ =Q^{\prime}}\ket{\psi}.$
Here, $C$ is a normalization constant and the following equation (true for any $q$ such that $q^{2}=\mathbbm{1}$, in particular $q\in\{X,Y,Z\}$) was used:
$\frac{\mathbbm{1}+sq}{2}q=s\frac{sq+\mathbbm{1}}{2}\quad\text{ for }s=\pm 1.$ (13)
Figure 2: Partial local measurement of graph states induces correlations in the remaining qubits. Pictures is the path graph $5$ vertices: if the associated graph state is measured according to the shown measurement pattern, Bell-like correlations are induced in the two terminal qubits.
As an example, consider the line graph of 5 qubits, as shown in Figure 2 and the Pauli string $O=X_{2}X_{3}X_{4}$ which has support on vertices in $I=\{2,3,4\}$. The stabilizer set of the 5 qubit line graph contains $2^{5}$ stabilizers. Three of them are $Q^{(a)}=X_{1}X_{3}X_{5}$, $Q^{(b)}=Z_{1}X_{2}X_{4}Z_{5}$, and $Q^{(c)}=-Y_{1}X_{2}X_{3}X_{4}Y_{5}$ which have support on qubits in $J^{(a)}=\{1,3,5\}$, $J^{(b)}=\{1,2,4,5\}$, and $J^{(c)}=\{1,2,3,4,5\}$, respectively. It holds that $Q^{(a)}|_{3}=O|_{3}$, $Q^{(b)}|_{2,4}=O|_{2,4}$, and $Q^{(c)}|_{2,3,4}=O|_{2,3,4}$. Following Lemma 1, the post-measurement state after measuring qubits 2, 3, 4 in the $X$-basis is stabilized by $s_{3}X_{1}X_{5}$, $s_{2}s_{4}Z_{1}Z_{5}$, and $-s_{2}s_{3}s_{4}Y_{1}Y_{5}$. This can be viewed as a graph state (Bell pair) in a nonstandard basis – the change is an effect of the measurement performed on the qubits $\{2,3,4\}$.
The remaining qubits are fully defined by the correlations developed during the measurement process. Thus, it is possible to find measurement schemes for which the output state has certain properties: e.g., if the goal is to extract a Bell pair from a larger graph state |$\ket{G}$, a measurement scheme $\prod_{i\in I}o_{i}$ must be found such that there exist stabilizers $Q$ of |$\ket{G}$ such that $Q_{I\cup J}=O_{I\cup J}$ and ensuring Bell-like correlations between the terminal vertices. For example in the 5 qubit line graph and the above chosen measurements, the Bell state stabilizers are extracted; more elaborate measurement patterns have been studied for use in quantum networks [20, 19, 27]. Further examples for graph families analyzed later in this article can be found in Appendix A.
As a direct result of the Lemma 1 we get the following observation:
Corollary 1.
If a stabilizer operator $Q=\pm\prod_{j\in J}q_{j}$ is fully embedded in the measurement scheme $O=\prod_{i\in I}o_{i}$ such that $J\subset I$ and $Q=\pm O|_{J}$, the only observable sign structures within $J$ are those such that $\prod_{j\in J}s_{j}=\pm 1$, where the sign is determined by the sign in $Q$.
The proof follows by considering only $J\subset I$ in Lemma 1.
As mentioned in the introduction, the creation of a Bell pair is equivalent to the task of information transfer (teleportation) across the graph. This is so because if the Bell pair creation in the graph $G$ is viewed in the context of a larger graph $G^{\prime}$, this process (followed by measurement of one of the terminal qubits) performs qubit fusion: the unmeasured terminal qubit acts as the two combined. Since all local measurements with nonoverlapping supports do commute, this can be done before any other measurement or afterwards: in the latter case, the information created in one of the qubits is merged with the second one. This is captured by the following Lemma, represented pictorially by the Figure 3.
Figure 3: If a measurement scheme generates a Bell pair across the vertices $\{{\operatorname{in}},{\operatorname{out}}\}$ in the state corresponding to a graph $G$, in a larger $G^{\prime}$ it can be used to transfer information across the graph. Extending the measurement by a properly chosen operator $A_{{\operatorname{in}}}$ leads to the fusion of the two terminal vertices.
Lemma 2.
Consider a graph $G=\{V,E\}$, and choose the two terminal vertices $\{{\operatorname{in}},{\operatorname{out}}\}$ such that $V=\{{\operatorname{in}},{\operatorname{out}}\}\cup I$. Suppose that the corresponding graph state |$\ket{G}$ undergoes a sequential measurement of $I$ producing a Bell pair across $\{{\operatorname{in}},{\operatorname{out}}\}$ in the meaning of Lemma 1.
With this determined, let us further assume that $G^{\prime}=(V^{\prime},E^{\prime})$ embeds $G$ in such a way that $\{{\operatorname{in}},{\operatorname{out}}\}$ is a separating set: the only connections between $G$ and the rest of $G^{\prime}$ are at these two vertices. Let us denote the vertices in $V^{\prime}\setminus V$ to which both qubits are connected by $\mathcal{N}^{\prime}({\operatorname{in}})$ and $\mathcal{N}^{\prime}({\operatorname{out}})$, respectively, and assume that the two sets are disjoint.
Within $G^{\prime}$, the aforementioned Bell pair-generating measurement followed by a product of single qubit unitary operations on the qubits $\{{\operatorname{in}},{\operatorname{out}}\}$ and the measurement $X_{\operatorname{in}}$ effectively performs qubit fusion of the terminal qubits. The new state is a graph state where the qubits $V^{\prime}\setminus V$ are intact, and the new neighborhood of the remaining qubit is $\mathcal{N}_{{\operatorname{out}}}\cup\mathcal{N}_{{\operatorname{in}}}$
Proof.
If within $G$ a Bell pair is produced across $\{{\operatorname{in}},{\operatorname{out}}\}$ as a result of the measurement pattern $O=\prod_{i\in L}o_{i}$ (where $L=V\setminus\{{\operatorname{in}},{\operatorname{out}}\}$), this means that there exist three $O$-consistent stabilizers involving the terminal vertices in the sense of Lemma 1. Combinatorical considerations show that one of those must be of a form such that the operators at the terminal vertices are $A_{\operatorname{in}},B_{\operatorname{out}}\in\{X,Y\}$ (but not $Z$), chosen independently.
Thus, there exists an $G$-stabilizer operator of the form ±$\pm A_{\operatorname{in}}B_{\operatorname{out}}\prod_{i\in I}o_{i}=g_{\operatorname{in}}g_{\operatorname{out}}\prod_{i\in I^{\prime}}g_{i}$, with $I,I^{\prime}\subset L$. Note that the definition of $g_{i}$ depends on the graph in question (Equation 6): if by $g^{\prime}_{i}$ we denote the generators corresponding to the entire graph $G^{\prime}$, the result is also a stabilizer, differing only by the $Z$ operators in the set $\mathcal{N}^{\prime}({\operatorname{in}})\cup\mathcal{N}^{\prime}({\operatorname{out}})$.
After the $A_{\operatorname{in}}$-basis measurement, we get a new stabilizer provided by Lemma 1:
(±$\left(\pm s_{\operatorname{in}}\prod_{i\in L}s_{i}\right)B_{\operatorname{out}}\mathchoice{}{}{}{}\prod_{j\in\mathcal{N}^{\prime}_{\operatorname{in}}\cup\mathcal{N}^{\prime}_{\operatorname{out}}}\mathchoice{}{}{}{}Z_{j},$ (14)
which can be taken to be a new stabilizer generator associated with the vertex ${\operatorname{out}}$, and brought to the canonical form of $X_{\operatorname{out}}\prod_{j\in\mathcal{N}^{\prime}_{\operatorname{in}}\cup\mathcal{N}^{\prime}_{\operatorname{out}}}Z_{j}$ by local unitary basis change of the ${\operatorname{out}}$-qubit. The properties of other stabilizers involving the qubits $\{{\operatorname{in}},{\operatorname{out}}\}$ can be proven similarly.
Thus, after the measurement within the Bell-generating part $L\cup\{{\operatorname{in}}\}$, the ${\operatorname{out}}$-qubit behaves exactly like it was connected to the neighborhood $\mathcal{N}^{\prime}_{\operatorname{in}}$ as well: the two vertices are fused and the information between them is transferred. ∎
4 Imperfect graph states
The reasoning presented above assumes that perfect $\operatorname{CZ}$ gates can be performed. However, this is not the case in experimental setups: some kind of noise is always present. In this section, we analyze three classes of physically motivated noise models; they all stem from different methods of graph state preparation. The origins and effects are sketched; for the detailed derivation, please refer to Appendix B.
Regardless of origin, noise processes often have a stabilizer-consistent description: vaguely speaking, they produce a mixture of stabilizer states, or the noise effect can be otherwise understood using the stabilizer-like correlations. In later sections, we show how to use this observation to mitigate the effects of noise.
4.1 Uncorrelated edge noise
Implementations like superconducting qubit quantum computers [28, 29], or ion traps [30, 31] prepare graph states by letting an initial product state evolve via an engineered Ising-like interaction Hamiltonian:
$H=\sum_{i\neq j}\gamma_{i,j}\frac{(Z_{i}+\operatorname{\mathds{1}})\otimes(Z_{j}+\operatorname{\mathds{1}})}{4}.$ (15)
Evolution according to such an interaction pattern effectively implements a sequence of controlled-phase operators $\operatorname{C\Phi}_{i,j}=\operatorname{diag}(1,1,1,\exp(-i\varphi_{i,j}))$. The structure of phases $\varphi_{i,j}=\gamma_{i,j}t$ is determined by the interaction strengths $\gamma_{i,j}$ and the evolution time $t$: inaccuracies in controlling either of them lead to the generation of a weighted graph state [22]. However, under general assumptions, the effective quantum channel implemented this way has a simple decomposition into local (one or two edges) operations, on top of the product of a sequence of $\operatorname{CZ}$ operations preparing the desired state.
Ideally, all of the generated phases $\varphi_{i,j}$ are equal to $\pi$. More realistically, with each experimental run, the actual realization will be $\varphi_{i,j}=\pi+\varepsilon_{i,j}$ with $\varepsilon_{i,j}\ll 1$. We discuss the two extremal cases of correlations of phase noises $\varepsilon_{i,j}$. The first, uncorrelated phase noise, assumes that the noises at different edges $(i,j)$ are completely independent. The other, correlated phase noise, assumes the contrary: each phase noise factor is equal to any other in a single experimental run, $\varepsilon_{i,j}=\varepsilon$. The noise encountered in experiments likely has yet a different structure with partial correlations between the edges; still, the two extremal processes may help in modeling the real-life scenario and strategies developed to mitigate them should apply in the more general cases.
Phases uncorrelated across the edges effectively lead to a probabilistic $\operatorname{CZ}$ (see the Appendix B) gate being implemented, if only each $\varepsilon_{i,j}$ is symmetrically distributed around 00 (so that $\varphi=\pi+\varepsilon$ is centered around $\pi$). Each of controlled-$Z$ unitary operators in Equation 1 is replaced with the quantum channel
$\operatorname{\mathcal{CZ}_{p}}[\rho]=(1-p)\operatorname{\mathcal{CZ}}[\rho]+p\rho,$ (16)
where $p$ depends on the distribution of $\varepsilon$ and $\operatorname{\mathcal{CZ}}[\rho]=\operatorname{CZ}\rho\operatorname{CZ}^{\dagger}$.
Thus effectively the effect of noise is the generation of randomized graph states [32]: ensemble of graph states built atop the original graph $G$ by removing the edge $(i,j)\in E$ with probability $p_{i,j}$; related graph ensembles have been introduced in [33] in a classical context. For the constant probability, the end state can be written as
$\rho=\sum_{E^{\prime}\subset E}(1-p)^{|E^{\prime}|}p^{|E\setminus E^{\prime}|}\ket{G^{\prime}}\bra{G^{\prime}},$ (17)
where $G^{\prime}=(V,E^{\prime})$ is a graph with edge set $E^{\prime}$ being a subset of the edge set $E$ of the original graph $G$.
Note that it is an effective description: in each experimental run, the prepared state is a pure weighted graph state [13]; the density operator in Equation 17 is a state of knowledge about the system, averaged over different noise realizations.
4.2 Correlated edge noise
If the weights in the Hamiltonian of Equation 15 are perfect ($\gamma_{i,j}=1$) but the interaction time is not perfectly controlled, within one experimental run, all the resulting phases $\varphi_{i,j}$ are equal. This is correlated phase noise: it has an effect similar, but more involved, to the uncorrelated case. Here, we also assume that $\varphi$ is symmetrically distributed around $\pi$. Let us assume that in each experimental run $\varphi_{i,j}=\varphi=\pi+\varepsilon$, and $\varepsilon$ is distributed with the normal distribution of the standard deviation $\sigma$. Then, with $p$ depending on the phase distribution (see Appendix B for details), the lowest order of approximation of the state (essentially, keeping only terms of order $\varepsilon^{2}$ in the end state and subsequent averaging) can be described with the help of square roots of the unitary operator $\operatorname{CZ}$:
$\rho=$ $(1-p\lvert E\rvert)\ket{G}\bra{G}+p\sum_{e\in E}\outerproduct{G-e}{G-e}$ | (18)
$+p\!\!\!\!\sum_{\{e,e^{\prime}\}\subset E\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \phantom{x}}^{e\neq e^{\prime}}\!\!\!\!\!\!\!\!\!\!\!\!\!\sum_{\phantom{x}\leavevmode\nobreak\ \leavevmode\nobreak\ \pm_{1},\pm_{2}}\!\!\!\!\!\!\!\pm_{1}\!\pm_{2}\!{\operatorname{\mathcal{CS}}_{e}^{\pm_{1}}}\!\!{\operatorname{\mathcal{CS}}_{e^{\prime}}^{\pm_{2}}}[\ket{G}\bra{G}].$
Note that this decomposition is not convex and cannot be interpreted as an ensemble: the minus signs in front of the decomposition terms prevent this. Signs appear as a result of discrete decomposition of controlled-phase and subsequent averaging; see Appendix B for details. In this decomposition, $G-e\coloneqq(V,E\setminus\{e\})$ is the graph with edge removed, and ${\operatorname{\mathcal{CS}}^{\pm}}$ is the application of one of the two unitary square roots of $\operatorname{CZ}$:
${\operatorname{\mathcal{CS}}^{\pm}}[\rho]=\sqrt{\operatorname{CZ}^{\pm}}\rho\sqrt{\operatorname{CZ}^{\pm}}^{\dagger},$ (19)
where
$\sqrt{\operatorname{CZ}^{\pm}}\coloneqq\frac{1}{2}\left[\left(1+e^{\pm i\frac{\pi}{2}}\right)\operatorname{\mathds{1}}+\left(1-e^{\pm i\frac{\pi}{2}}\right)\operatorname{CZ}\right].$ (20)
4.3 Local $Z$ flip noise
In linear optic experiments, where Bell pairs are generated and subsequently fused to get a graph state, the entangling operation may fail as a result of partial photon distinguishability (by imperfect frequency of spatial mode overlap) [34, 35, 36] – for a derivation see the Appendix B. The effect of a noisy fusion gate described this way can be modeled as a perfect fusion followed by a probabilistic application of the $Z$ unitary to the surviving optical qubit $i$:
$F_{i}(\rho)=(1-p)\rho+pZ_{i}\rho Z_{i}.$ (21)
Here, the probability $p$ depends on the level of photon distinguishability of the fused qubits. Thus, the final graph state developed in this procedure can be thought of as a probabilistic application of local $Z$ unitaries to each of the qubits in the graph.
4.4 Quantification of the noise effects
In the following chapters, we determine the entanglement quality of a small (two- or three-qubit) subsystem left after a measurement procedure performed on a graph state: this is the Bell pair or GHZ state extraction, since ideally the resulting state would represent correlations found in these states. We have chosen fidelity with respect to the ideal state as a measure of the entanglement quality. In the absence of noise, the initial state |$\ket{G}$ is pure, and measurement of an $i$-th qubit in the basis of an operator $o_{i}\in\{X_{i},Y_{i},Z_{i}\}$ leads to observation of an outcome ±$\pm 1$. On the algebraic level, this corresponds to application of an operator $\frac{1}{2}\left(\mathbbm{1}\pm o_{i}\right)$ to the pre-measurement state, and after the entire procedure, the post-measurement state |$\ket{\psi}$ is still pure. This allows a simple determination of the fidelity with respect to the noisy state $\rho$: if the post-measurement state is $\sigma$, the following holds.
(|$\mathcal{F}(\ket{\psi}\!\bra{\psi},\sigma)=\langle\psi|\sigma|\psi\rangle.$ (22)
Note that the ideal post-measurement state depends on the observed outcomes during the experiment: they do affect the signs in the final stabilizer structure (see Lemma 1). Therefore, the ideal state |$\ket{\psi}$ must be determined from the signs of the outcomes. This is only possible if the structure is consistent with the stabilizers of the ideal state; this is not an issue, since we explicitly postselect on the right set of outcomes in the following sections as part of the error mitigation strategy. Finally, the average fidelity $\langle\mathcal{F}\rangle$ is determined for all possible measurement outcomes consistent with the stabilizer structure; this is effectively approximated by taking $N$ samples from the defining ensembles for a specific noise model with a Monte Carlo algorithm and thus
$\langle\mathcal{F}\rangle=\frac{1}{N}\!\!\!\sum_{\text{outcomes}}\!\!\!\mathcal{F}(\text{outcome}).$ (23)
Since we are mostly interested in the behavior in the low noise limit, we define fidelity susceptibility $\alpha$ as the rate of change of mean fidelity $\langle\mathcal{F}\rangle(p)$: the ensemble of extracted and postselected states depends on the noise, and the susceptibility captures the effects of small but not negligible noise.
[…]
For zero noise, the fidelity is equal to 1 by definition. For any other amount, the fidelity may only decrease, and hence the susceptibility is positive. Its numerical magnitude determines how fragile the extraction procedure is to the effects of noise: a robust one has a small (or, in some cases, even zero) susceptibility.
5 Extraction from noisy ensembles
In this section we apply the methods developed in previous parts to show the effects of the noise and find measurement patterns that minimize them. Instead of focusing on teleportation directly, we analyze the (equivalent) task of extracting a high-quality Bell pair from a larger noisy state, close to a graph state |$\ket{G}$.
5.1 Noise-correcting structures and robust families of states
As introduced above, noise processes lead to an ensemble of different states $\{(p_{i},\ket{\psi_{i}}\}$ being prepared, where one of the ensemble components is the ideal state |$\ket{G}$. This ensemble can be interpreted as a probabilistic preparation of an unknown state |$\ket{\psi_{i}}$ with probability $p_{i}$. The same measurement procedure as previously might now yield results inconsistent with |$\ket{G}$, and observations of those are an indication of state other than |$\ket{G}$ being prepared.
As an example, consider the graph $G$ to be a $4$-cycle of vertices $\{1,2,3,4\}$, as shown in Figure 1a. The operator $O=X_{1}X_{3}$ belongs to a stabilizer of |$\ket{G}$, thus the only sign structures that can be observed in the sequential measurement of qubits $1$ and $3$ are $+1,+1$ and $-1,-1$. Consider now all edge subgraphs $\{G^{\prime}\}$ of $G$ and a physical process in which an ensemble of graph states $\{p_{G^{\prime}},\ket{G^{\prime}}\}$ is prepared, some of them are shown in Figure 1b. The operator $O$ (and $-O$) is not a part of the stabilizer of any proper subgraph with at least one edge; thus, sign structures of $+1,-1$ or $-1,+1$ are possible: observing those in the same measurement procedure implies that the pre-measurement state was not |$\ket{G}$.
Figure 4: Considered families of graphs for the task of Bell pair extraction, parameterized by the length of the internal part $n$. The terminal qubits, designated for creation of a Bell pair, are denotes with filled nodes. From the top: path graph, twisted pair, crazy graph.
The postselection on the right measurement outcomes consistent thus performs a probabilistic check whether |$\ket{G}$ was truly prepared. However, the existence of such parity-checking structures depends on the graph and the task for which the associated quantum state is used.
A Bell pair can be prepared with the help of a path graph state and local $X$ measurements, but such a system does not support embedded stabilizer operators. Other graphs and measurement patterns do better in this regard, and here we present them. The families of graph states analyzed are parameterized by the length between the terminal qubits across which a Bell pair is to be prepared (see Figure 4). Apart from a simple path graph, included as a benchmark for comparison, we include two other families described below – both of them have their strengths.
The twisted pair graph [37] is built from layers, where each qubit in the layer $k$ is connected with every qubit from the layers $k-1$ and $k+1)$. The number of qubits in each layer alternates between $1$ and $2$, which enables the existence of a restricted set of stabilizers: for each 2-qubit layer consisting of the qubits $\{i,j\}$, the operators $X_{i}X_{j}$ stabilize |$\ket{G}$. Additionally, it is very robust to certain types of noise, because of its relatively simple geometry.
The crazy graph (studied in [38, 14] for its noise robustness) also consists of layers of $2$ qubits each, with a similar full connectivity structure between the adjacent layers. This ensures a robust and simple to analyze structure of the embedded stabilizer operators (see Appendix A and Figure 5).
Figure 5: Crazy graph of length 3 together with an all-$X$ measurement pattern possess embedded error-checking and correction structures. Middle: each layer is associated with an $X\otimes X$ stabilizer, which can be used to confirm the proper preparation of the state. Bottom: even if any single internal edge (dotted line between layers 1 and 2) is missing, the terminal qubits correlations are not affected: the operators shown are stabilizers, regardless of whether the edge is present or not.
Both presented structures generalize to the extraction of other types of states, in which context they serve as a direct replacement of path graphs. We mention the usage in a task of 3-qubit GHZ state extraction.
5.2 Bell pair quality scaling for different type of noise
Each of the graph states associated with the aforementioned three families (path graph, twist graph, and crazy graph) can be used for extraction of a Bell pair across the terminal qubits. Only local $X$ measurements are used for all three, and the fact that in the ideal case a maximally entangled state is produced is captured by the stabilizer operators consistent with this measurement pattern.
As a simple example, consider the 5 qubit line graph as discussed in Section 3. Measurements in the $X$ basis on all but the terminal qubits lead to a two-qubit state which can be shown to have Bell-like correlations. Similar structures can be found for all the three graph families mentioned: sets of stabilizer operators ensuring correlations of the terminal qubits, having only $X$ operators in the internal section (and thus compatible with the measurement scheme).
The effect of noise destroys the perfect correlations: even if for every ensemble element the end state is maximally entangled, the mixture of such states might have less (or none) entanglement. Postselection on the measurement outcomes consistent with the ideal stabilizer structure helps twofold: one of its effects is probabilistic error detection. The other, appearing in the case of the crazy graph ensemble, is more subtle: after postselection, the end state may be the same for a large portion of the ensemble elements (see Appendix A).
The results, presented in Figure 6, show that this approach (noise-detecting structures combined with postselection) works for different types of noise. The fidelity susceptibility (quantifying the initial fidelity falloff in the low noise limit; see Equation 24) is reduced for the twisted pair and crazy structures. Furthermore, the susceptibility for the crazy graph structure does not depend on the length, as a result of the correctional stabilizers mentioned in the previous paragraph. This behavior is observed both for uncorrelated phase noise, resulting in probabilistic edge losses and a local $Z$ flip noise.
An additional entanglement-preserving structure can be found in the case of perfectly correlated phase noise. The authors of Ref. [13] recently observed that in the case of the path graph, postselection on different measurement outcomes yields vastly different results in the terms of entanglement quality; we have found that this result does generalize to the other linear graphs analyzed by us. Both for the crazy and double twist graph families the lowest order nontrivial noise effects can be canceled completely if all the measurement outcomes are postselected on measuring the minus outcome of $X$. This is consistent with our previous observations: such an observed measurement pattern does not violate parity constraints arising from the stabilizer structure. This result can be understood by studying the modifications of the stabilizer operators of |$\ket{G}$ under correlated controlled-phase noise.
For the square graph presented in Figure 1a, if instead of the perfect controlled-$Z$ unitary, the $\operatorname{C\Phi}(\pi+\varepsilon)=\operatorname{diag}(1,1,1,-1\exp(i\varepsilon))$ is used (so that $\operatorname{C\Phi}=\operatorname{CZ}$ for $\varepsilon=0$), the four-qubit entangled state is a weighted graph state |$\ket{G}_{\varepsilon}$. If qubits 1 and 3 are measured in the $X$ basis, we expect the measurement outcomes to be equal (we exclude the cases when it does not hold). With the weighted graph state |$\ket{G}_{\varepsilon}$ as the initial state, let us denote the unnormalized post-measurement state by
|$\ket{X_{1,3}\pm}=\frac{\mathbbm{1}\pm X_{1}}{2}\frac{\mathbbm{1}\pm X_{3}}{2}\ket{G}_{\varepsilon}.$ (25)
The relevant expectation values defining the end correlations (with $\langle O\rangle_{\pm}$ denoting the expectation value of $O$ for the unnormalized state |$\ket{X_{1,3}\pm}$) across qubits $2$ and $4$ are the following, in the lowest nontrivial noise order:
$\langle Z_{2}Z_{4}\rangle_{+}$ $=1-\varepsilon^{2},$ $\langle Z_{2}Z_{4}\rangle_{-}$ $=-1+\frac{\varepsilon^{2}}{2},$ (26)
$\langle X_{2}X_{4}\rangle_{+}$ $=1-3\varepsilon^{2},$ $\langle X_{2}X_{4}\rangle_{-}$ $=-1+\frac{\varepsilon^{2}}{2},$
$\langle Y_{2}Y_{4}\rangle_{+}$ $=-1+3\varepsilon^{2},$ $\langle Y_{2}Y_{4}\rangle_{-}$ $=1+\frac{\varepsilon^{2}}{2},$
$\||X_{1,3}+\rangle\|^{2}$ $=1-\varepsilon^{2},$ $\||X_{1,3}-\rangle\|^{2}$ $=1-\frac{\varepsilon^{2}}{2}.$
Note that for the |$\ket{X_{1,3}-}$ outcome the squared length $\||X_{1,3}-\rangle\|^{2}$ is proportional to the unnormalized expectation values: thus, for the physical correlations the dominant noise effects are canceled and only higher order terms ($O(\varepsilon^{4})$) remain. On the other hand, if the outcome is |$\ket{X_{1,3}+}$, the noise is amplified compared to postselection on only stabilizer-consistent outcomes. Thus, further postselection can amplify the entanglement quality of the remaining qubits at the cost of discarding certain outcomes and reduced production rates.
The double twist graph can be viewed as multiple such squares stacked together by the terminal vertices: thus, a proper postselection can mitigate the first nontrivial noise effects completely (see Figure 7). A similar structure appears for the crazy graph: it is much more complex to analyze it, but computer algebra systems provide evidence of a periodic structure of susceptibilities as the length of the graph increases.
Figure 6: Susceptibility of the Bell pair entanglement quality to edge loss (top) and local phase flip (bottom) noise. Pictured is the linear part of concurrence decrease for the three analyzed families (with postselection on the measurement outcomes consistent with the ideal state) in the low noise limit, where the dominant contributions are the single edge losses (top) and single qubit flips (bottom). The constant susceptibility of the crazy graph is a result of the noise at the terminal edges only: each internal edge loss is automatically corrected for and every single flipped qubit is detected.
Figure 7: Susceptibility of the Bell pair entanglement quality for correlated phase noise. Here, only the cases postselected on observing exclusively minus signs during the $X$ measurement are taken into account. In this regime, the twist graph is able to remove the effect of noise completely (in the lowest nontrivial orders), and the crazy graph shows a periodic pattern of susceptibilities, still significantly reducing the noise effect. The situation inverts for observing exclusively $+$ signs: then, the noise gets amplified.
6 Extraction of GHZ states
The methodology applied for Bell pair extraction generalizes to the extraction of GHZ states [23]: generally, for $N$ qubits, they correspond to a star graph with a central qubit connected to the remaining $N-1$ ones (or, equivalently, a complete graph of $N$ qubits). Here, we concentrate on the case $N=3$ and adopt star-shaped graph structures built from the graph patterns used in Bell pair creation. Refer to Figure 8a for an illustration of the structure: it is built on the template of several crazy graphs, stitched at the central pair of vertices. The fidelity response, presented in Figure 8b, is more resilient to noise (as compared to the simpler path graph) also in this new task of extraction of the GHZ state.
The data presented in this figure require calculations involving a large number of ensemble states and measurement outcomes. To mitigate the numerical difficulty, a Monte Carlo method was utilized. The figure of merit – fidelity – is linear with respect to the ensemble- and measurement outcome averaging, provided the target state is pure. Thus, this approach allows for an approximation of fidelity by sampling ensemble elements and measurement outcomes, avoiding direct calculations with density operators. The fidelity of the output state to the ideal pure state is estimated using this method until the plots converge smoothly and the variance of the fidelity estimator is negligible.
Analysis of the data in Figure 8 shows that the fidelity response for both the GHZ and Bell state extractions exhibits qualitatively similar behavior: most importantly, the susceptibility is greatly reduced in the low noise limit. Thus, the ’crazy graph’ structures can generalize beyond Bell pair extraction to more complex multi-qubit systems, maintaining their noise-resilient properties. This performance gain is due to the noise-detecting and -correcting capabilities intrinsic to these graph families. Further investigations of different connectivity structures suggest that this behavior is universal, although a formal proof is beyond the scope of this study. If true, the ’crazy graph’ can serve as a link similar to a simple path link, similar in function but implementing some level of error correction.
a)Refer to caption
b)Refer to caption
Figure 8: a) Structures of the analyzed graphs analyzed: path and crazy templates yielding Bell pairs and 3-qubit GHZ states upon measurements of the central qubits. b) Fidelity of the end state with respect to the ideal Bell and GHZ states as a function of noise. The fidelity is averaged over possible measurement outcomes consistent with the measurement-embedded stabilizer structures. The initial fall-off rate of fidelity at $p=0$ is the fidelity susceptibility, defined in Equation 24.
7 Summary and Outlook
This study has advanced the understanding of noise processes in graph states, specifically addressing the effects of preparation by the Ising-like interaction and photonic qubit fusion. We have found graph structures applicable in error detection and correction, especially in tasks such as information transfer and GHZ state preparation. This approach uses additional qubits for probabilistic verification and demonstrates a new method to improve resilience against specific noise disturbances.
Future research may include more general tasks found in the measurement-based quantum computation approach. So far, our results are applicable to measurement patterns composed of Pauli operators, but it is known to restrict the set of possible operations to a clasically simulable one: a similar method, based on non-Pauli measurements, but still allowing for noise effects reduction, could open new pathways for quantum computation, albeit with conceptual and practical complexities. The methods presented here can be generalized for arbitrary output graph geometry, and the applicability of this procedure for quantum metrology [39] could be studied.
Additional platforms for generating graph states can also be analyzed: in this context, sequential single-atom emitters [25] are especially interesting. Although noise processes are fairly complex, such systems are capable of producing extensive path graph states and present a unique opportunity for advancing graph state generation techniques. Different experimental realizations natively support different connectivity structures [24, 25, 30, 40, 41], and optimal utilization of a given experimental setting may require further development of the presented techniques.
To summarize, this research contributes new insights into noise reduction in graph state-based systems, implications of which include applications of measurement-based quantum computation. We hope to apply a similar approach for different applications and determine its utility as graph states are realized on existing quantum devices.
Acknowledgments
We thank Frederik Hahn, Mariami Gachechiladze, and Jan L. Bönsel for discussions. This work was supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation, project numbers 447948357 and 440958198), the Sino-German Center for Research Promotion (Project M-0294), the ERC (Consolidator Grant 683107/TempoQ), the German Ministry of Education and Research (Project QuKuK, BMBF Grant No. 16KIS1618K), the Stiftung der Deutschen Wirtschaft, and the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement (No 945422).
Appendix A Post-measurement correlations and stabilizers compatible with the measurement scheme
Lemma 3.
Given a graph state |$\ket{G}$ of $N$ qubits and a certain measurement scheme defined in Lemma 1 by a set of qubits $I$ and operators $o_{i}$, then the subset of the stabilizer of |$\ket{G}$ compatible with the measurement pattern has $N-\lvert I\rvert$ generators, and therefore fully defines the post-measurement state.
Proof.
The stabilizer of |$\ket{G}$, up to the structure of the signs, can be viewed as a linear subspace of $\mathbb{Z}_{2}^{2N}$ – this is known as the binary representation of the stabilizer [5]. The identification is following: if the stabilizer operator is defined as $\prod_{j\in J}g_{j}$, the corresponding element of $\mathbb{Z}_{2}^{2N}$ is $(x_{1},\ldots,x_{N})\oplus(z_{1},\ldots,z_{N})$, with $x_{i}$ being 1 if and only if $i\in J$ and $\vec{z}=A_{G}\vec{x}\mod 2$, where $A_{G}$ is the adjacency matrix of $G$. Thus, if a stabilizer contains $X_{i}$, the corresponding numbers are $(x_{i},z_{i})=(1,0)$, for $Z_{i}$ it is $(x_{i},z_{i})=(0,1)$, and $Y_{i}$ corresponds to $(x_{i},z_{i})=(1,1)$; no operator at site $i$ is denoted by $(x_{i},z_{i})=(0,0)$. In this way, the additive algebra of $\mathbb{Z}_{2}^{2}$ mimics the product rules of Pauli operators with only the missing sign structure.
The stabilizer of |$\ket{G}$ corresponds to a special $N$-dimensional subspace in $\mathbb{Z}_{2}^{2N}$ [5]. For any stabilizer operator to be compatible with the measurement scheme in the sense of Lemma 1, each operator of its Pauli string form with support in $I$ must be $o_{i}$: this introduces constraints to the set of compatible operators. However, these constraints translated into the language of $\mathbb{Z}_{2}^{2N}$ are linear: if $X_{i}$ is measured at site $i$, the additional constraint is $z_{i}=0$ (this encompasses the cases of $X_{i}$ and identity at site $i$), if $Z_{i}$ is measured, the constraint is $x_{i}=0$, and for $Y_{i}$ it is $x_{i}+z_{i}=0\mod 2$. All of them are independent, and there are $\lvert I\rvert$ of them, one for each measured site. Thus, the linear subspace of $\mathbb{Z}_{2}^{2N}$ corresponding to measurement-compatible stabilizers has dimension $N-\lvert I\rvert$; any basis of this subspace corresponds to generators of the post-measurement state stabilizer. ∎
For example, the correlations that stem from the measurement described after Lemma 1 and in Figure 2 correspond to the following $\vec{x}\oplus\vec{z}$ vectors in the binary representation:
$(0,1,0,1,0)$ $\oplus(1,0,0,0,1),$ (27)
$(1,0,1,0,1)$ $\oplus(0,0,0,0,0),$
$(1,1,1,1,1)$ $\oplus(1,0,0,0,1).$
The third vector is a linear combination of the first two; thus the space of end correlations is two-dimensional.
A.1 Patterns for specific graph families
Consider a path graph $G$ of $n+2$ vertices $V=\{0,1,\ldots,n,n+1\}$ with edges $E=\{\{k,k+1\}\}_{k=0}^{n}$. If the qubits $\{1,\ldots,n\}$ of the associated state |$\ket{G}$ are measured in the basis of $X$, this induces correlations in the remaining unmeasured qubits 00 and $n+1$. This follows from the fact that the following stabilizer operators are consistent with the measurement pattern:
$Z_{0}Z_{n+1}\prod_{k\in I}X_{k},$ $\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ X_{0}X_{n+1}\prod_{k\in I^{\prime}}X_{k},$ (odd $(\text{odd }n)$ (28)
$Z_{0}X_{n+1}\prod_{k\in I}X_{k},$ $\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ X_{0}Z_{n+1}\prod_{k\in I^{\prime}}X_{k}.$ (even $(\text{even }n)$
where $I$ ($I^{\prime}$) is the set of odd (even) numbers between $1$ and $n$:
$I$ =$=\{k\in\{1,2,\ldots,n\}:k=1\mod 2\},$ (29)
$I^{\prime}$ =$=\{k\in\{1,2,\ldots,n\}:k=0\mod 2\}.$
Thus, if the outcome of the $X_{k}$ measurement is denoted by $s_{k}$, the following operators generate the post-measurement stabilizer of the terminal qubits $\{0,n+1\}$:
$Z_{0}Z_{n+1}\prod_{k\in I}s_{k},$ $\leavevmode\nobreak\ X_{0}X_{n+1}\prod_{k\in I^{\prime}}s_{k},$ (odd $(\text{odd }n)$ (30)
$Z_{0}X_{n+1}\prod_{k\in I}s_{k},$ $\leavevmode\nobreak\ X_{0}Z_{n+1}\prod_{k\in I^{\prime}}s_{k}.$ (even $(\text{even }n)$
Both cases are consistent with a maximally entangled two-qubit state being produced in the terminal qubits: for the odd-$n$ case, the existence of stabilizers described by Equation 30 ensures that further measurement of $X$, $Y$ or $Z$ at qubit 00 is perfectly (anti)correlated with the outcome of measurement of the same operator at qubit $n+1$.
This result generalizes to the case where each qubit $k$ is replaced by a set of $m_{k}$ qubits in such a way that the adjacent layers are fully connected. In this case, the qubits are denoted by pairs of numbers $(k,i)$, where $k=\{0,1,\ldots,n,n+1\}$ is the layer index and $i\in\{1,\ldots,m_{k}\}$ is the qubit index within the layer. The edges exist between any layer-adjacent qubit:
$E=\{\{(k,i),(k+1,j)\}:k\in\{0,\ldots,n\},i\in\{1,\ldots,m_{k}\},j\in\{1,\ldots,m_{k+1}\}\}.$ (31)
This class of graphs includes the path graph (each layer has only a single qubit), the crazy graph (the layers 00 and $n+1$ contain one qubit each, every other has two) and twisted graphs (the qubit count alternates between 1 and 2) as well. Suppose that all the internal qubits in every layer are measured in the $X$ basis. The stabilizer operators that determine the remaining correlations must be consistent with this structure. A particularly simple set of such operators exists: first, observe that if a layer $k$ contains at least two qubits, $X_{k,i}X_{k,i^{\prime}}$ is a stabilizer operator. Thus, the signs of outcomes have to be equal within a layer. Therefore, for any choice of inner-layer qubit index $(i_{k})$ the direct analogue of Equation 28 holds:
$Z_{0}Z_{n+1}\prod_{k\in I}X_{k,i_{k}},$ $\leavevmode\nobreak\ X_{0}X_{n+1}\prod_{k\in I^{\prime}}X_{k,i_{k}},$ (odd $(\text{odd }n),$ (32)
$Z_{0}X_{n+1}\prod_{k\in I}X_{k,i_{k}},$ $\leavevmode\nobreak\ X_{0}Z_{n+1}\prod_{k\in I^{\prime}}X_{k,i_{k}},$ (even $(\text{even }n),$
are stabilizer operators. For symmetry, we assume the twisted graph structure exists only for an odd number of layers, so that the terminal ones consist only of one qubit each. Thus, the terminal layers 00 and $n+1$ always consist only of one qubit each, and the inner indices are omitted. Consequently, if $s_{k}$ in Equation 30 are interpreted as signs of outcomes at layer $k$ (which are all equal within a layer due to the structure of the stabilizer operators mentioned previously), the same post-measurement stabilizer structure appears.
In the case of the crazy graph, an additional observation can be made: loss of any single internal (not adjacent to the terminal qubits) edge yields the same post-measurement stabilizer structure, since the indices $(i_{k})$ in Equation 32 can be chosen arbitrarily. Thus, the crazy graph family is especially tolerant to single edge loss: it does not affect the post-measurement state, if the outcomes are postselected on the ideal sign structure. This can be generalized to the loss of any subset of edges not adjacent to each other and the terminal qubits.
Appendix B Derivation for the noise models
B.1 Uncorrelated phase noise
Figure 9: The controlled-$Z$ gate, if implemented as controlled phase operation, is susceptible to phase variations. In each experimental run the phase is constant, and the resulting state is a weighted graph state. Since this phase is unknown, the effective quantum state is weighted graph state, averaged over different noise realization. For just a single edge, if the phase distribution is centered around $\pi$ (corresponding to the ideal graph state), the resulting quantum channel is a mixture of controlled-$Z$ and identity: the edge is probabilistically generated.
In certain ion traps and superconducting quantum devices, the sequence of $\operatorname{CZ}$ gates is realized with evolution with an Ising-type Hamiltonian:
$H=\sum_{(i,j)\in G}\gamma_{i,j}\frac{Z_{i}+1}{2}\frac{Z_{j}+1}{2}.$ (33)
For constant $\gamma=\gamma_{i,j}$, the unitary operation $\exp(-iHt)$ exactly corresponds to the product of $\operatorname{CZ}$ gates appearing in Equation 1 for $t=\frac{\pi}{\gamma}$. However, incomplete control over the system leads to a fluctuation of the interaction strengths $\gamma_{i,j}$. In such case, each of the two-body terms appearing in Equation 33 gives rise to a controlled phase operation $\operatorname{C\Phi}_{i,j}=\operatorname{diag}(1,1,1,\exp(-it\gamma_{i,j}))$ in the appropriate two-body subsystem. The quantum channel corresponding to this gate decomposes:
$\operatorname{C\Phi}\rho\operatorname{C\Phi}^{\dagger}=\frac{1-\cos\phi}{2}\operatorname{\mathcal{CZ}}[\rho]+\frac{1+\cos\phi}{2}\rho+\frac{\sin\phi}{2}\operatorname{\mathcal{CS}}^{+}[\rho]-\frac{\sin\phi}{2}\operatorname{\mathcal{CS}}^{-}[\rho],$ (34)
where $\phi=t\gamma_{e}$ and $\operatorname{\mathcal{CS}}^{\pm}$ is defined in Equation 19. If the actually realized phase cannot be measured classically in each experimental run, the resulting state is a mixture of states:
[…]
where $P(\{\gamma_{e}\})$ denotes the probability distribution of phases and |$\ket{G_{\gamma}}=\left(\prod_{i,j}\operatorname{C\Phi}_{i,j}\right)\ket{+}^{\otimes N}$. If the strength noise is independent for each edge (i.e. $P=\prod P_{e}$) and symmetric around $\pi$, a significantly easier description of the state $\rho$ can be found: each controlled-phase averages to a mixture of controlled-$Z$ and identity (see Figure 9), since these are the only terms appearing in Equation 34 with prefactors symmetric around $\pi$. Thus, noise effectively leads to the generation of randomized graph states: this is an ensemble of graph states, where the edge $e\in E$ is missing with probability $p_{e}$:
$\rho=\sum_{E^{\prime}\subset E}\left(\prod_{{\{i^{\prime},j^{\prime}\}\in E^{\prime}}}\!\!\!(1-p_{i^{\prime},j^{\prime}})\!\!\!\prod_{\{i,j\}\notin E^{\prime}}\!\!\!p_{i,j}\right)\ket{G^{\prime}}\bra{G^{\prime}},$ (36)
where the graph $G^{\prime}$ has the edges defined by the set $E^{\prime}$. The probability $p_{i,j}$ can be determined by the statistical properties of the distribution of the phase $\phi$ associated with the edge $\{i,j\}$. In particular, for the normal distribution of $\varepsilon$ centered around $\pi$ with the standard deviation of $\sigma$ the probability $p$ takes the closed form of
$p=\frac{1-\exp\left(-\frac{\sigma^{2}}{2}\right)}{2}=\frac{\sigma^{2}}{4}-\frac{\sigma^{4}}{16}+O(\sigma^{6}).$ (37)
B.2 Correlated phase noise
Correlated phase noise has similar, but more involved effect. Suppose that the weights in Hamiltonian Equation 33 are all the same, but the interaction time varies. In such a case, the phase is the same for each controlled-phase operator, and the decomposition used to derive Equation 17 breaks. This is due to the fact that in the integrand of Equation 35 there exist additional terms symmetric around the ideal phase of $\phi=\pi$. There are the terms proportional to $\sin\phi$ in Equation 34: even powers of them also contribute to the final state. Thus, in the lowest order or approximation the output state is described by:
  • The unmodified state graph state |$\ket{G}$ with reduced probability coming from $\operatorname{\mathcal{CZ}}$ terms appearing in product of channels defined by Equation 34,
  • graph states with one edge missing as in the case of uncorrelated noise,
  • graph states modified by a product of two $\operatorname{\mathcal{CS}}^{\pm}$ channels corresponding to different edges, with possible negative weights coming from negative signs in Equation 34.
In the lowest order of approximation, corresponding to the $\sigma^{2}$ terms in Equation 37, the calculation yields the final state (after averaging) of the form
$\rho=$ $|\displaystyle(1-p\lvert E\rvert)\ket{G}\bra{G}+p\sum_{e\in E}\lvert{\!\!\overbrace{\!G^{\prime}\!}^{G\setminus\{e\}}\!\!}\rangle\bra{G^{\prime}}$ | (38)
$+\frac{p}{2}\sum_{(e,e^{\prime}\in E}^{e\neq e^{\prime}}\sum_{\pm_{1},\pm_{2}}\pm_{1}\pm_{2}\operatorname{\mathcal{CS}}_{e}^{\pm_{1}}\operatorname{\mathcal{CS}}_{e^{\prime}}^{\pm_{2}}[\ket{G}\bra{G}].$
Here, $\lvert E\rvert$ denotes the number of edges in the graph $G$. Note that it is not an ensemble decomposition: in the last sum, some signs are negative. This decomposition is not unique, and in principle different weights for the full, edge-missing, and $\operatorname{\mathcal{CS}}$-modified graph states are possible; however, some of the signs will always be negative, regardless of the choice. However, such decompositions are only low noise approximations for $p\ll\lvert E\rvert^{-1}$.
B.3 Fusion gate noise
Figure 10: Polarizing beamsplitter transmits horizontally polarized photons and reflects vertically polarized ones, entangling the input state in the output modes. Subsequent polarization measurement in the diagonal basis of one of the output modes is a basis for type 1 fusion gate, which is used in generation of larger graph states from Bell pairs.
Optical realizations [41], while using different mechanisms for entangling, lead to a similar result. If a type 1 entangling gate is used, imperfect mode matching leads affected qubits to be effectively depolarized in the $Z$ basis. This has the exact same result on the output state as local (single-qubit) phase noise, despite different physical origins. A different scheme for generating large cluster states is employed in linear optics experiments. In such systems, the qubits are encoded as photons in different light modes. Two modes are associated with each qubit, and a single photon being in one of them corresponds to the standard computational basis states. Usually, these are polarization (horizontal/vertical) modes with overlapping spatial profiles. Thus, let us denote the creation operators of those modes by ${\hat{h}^{\dagger}}_{i}$ and ${\hat{v}^{\dagger}}_{i}$.
In this context, a Bell pair encoded as the nontrivial two-vertex graph state can be written as
|$\ket{{\bullet\!\!-\!\!\bullet}}=\frac{1}{2}(1,1,1,-1)^{T}=\frac{{\hat{h}^{\dagger}}_{1}{\hat{h}^{\dagger}}_{2}+{\hat{v}^{\dagger}}_{1}{\hat{h}^{\dagger}}_{2}+{\hat{h}^{\dagger}}_{1}{\hat{v}^{\dagger}}_{2}-{\hat{v}^{\dagger}}_{1}{\hat{v}^{\dagger}}_{2}}{2}\ket{\emptyset},$ (39)
where $\ket{\emptyset}$ denotes the vacuum state of the optical field. For creation of a larger graph states, $N$ such Bell pairs generated through spontaneous parametric downconversion, with the initial state of
|$\ket{{\bullet\!\!-\!\!\bullet}}^{\otimes N}=\prod_{i=1}^{N}\frac{{\hat{h}^{\dagger}}_{2i-1}{\hat{h}^{\dagger}}_{2i}+{\hat{v}^{\dagger}}_{2i-1}{\hat{h}^{\dagger}}_{2i}+{\hat{h}^{\dagger}}_{2i-1}{\hat{v}^{\dagger}}_{2i}-{\hat{v}^{\dagger}}_{2i-1}{\hat{v}^{\dagger}}_{2i}}{2}\ket{\emptyset}.$ (40)
The Bell pairs are subsequently fused to create a larger graph state. Every fusion takes two spatial modes (indexed by numbers in the subscription) and mixes them through a polarizing beamsplitter – see Figure 10 for the details. Then, one of the output modes is measured in the diagonal $h\pm v$ polarization basis. To see the effect of such operation, consider arbitrary input graph state (not necessarily a tensor product of Bell pairs) for graph $G=(V,E)$ such that the vertices 1 and 2 are not connected:
|$\ket{G}$ =$=\prod_{\{1,j\}\in E}\operatorname{CZ}_{1,i}\prod_{\{2,j\}\in E}\operatorname{CZ}_{2,j}\ket{+}\otimes\ket{+}\otimes\ket{G^{\prime}}$ (41)
$=\frac{{\hat{h}^{\dagger}}_{1}{\hat{h}^{\dagger}}_{2}f_{++}+{\hat{v}^{\dagger}}_{1}{\hat{h}^{\dagger}}_{2}f_{-+}+{\hat{h}^{\dagger}}_{1}{\hat{v}^{\dagger}}_{2}f_{+-}+{\hat{v}^{\dagger}}_{1}{\hat{v}^{\dagger}}_{2}f_{--}}{2}\ket{\emptyset},$
where the symbols $f$ denote expressions involving creation operators to arrive at the state |$\ket{\psi}$. The $f_{++}$ prepares the graph state |$\ket{G^{\prime}}$, while $f_{-+}$ creates the state $\prod_{(1,i)\in G}z_{i}\ket{G^{\prime}}$: it is the postselected graph state upon measuring negative sign of the first qubit in the $Z$ basis. The other two operators work similarly; however, $f_{--}$ flips the phase of |$\ket{G^{\prime}}$ only in the qubits that are connected to only $1$ or $2$.
Polarizing beamsplitter exchanges the vertical polarization, keeping the horizontal ones intact ($U{\hat{h}^{\dagger}}_{1,2}={\hat{h}^{\dagger}}_{1,2}U$, $U{\hat{v}^{\dagger}}_{1,2}={\hat{v}^{\dagger}}_{2,1}U$):
$U\ket{G}$ $=\frac{{\hat{h}^{\dagger}}_{1}{\hat{h}^{\dagger}}_{2}f_{++}+{\hat{v}^{\dagger}}_{2}{\hat{h}^{\dagger}}_{2}f_{-+}+{\hat{h}^{\dagger}}_{1}{\hat{v}^{\dagger}}_{1}f_{+-}+{\hat{v}^{\dagger}}_{1}{\hat{v}^{\dagger}}_{2}f_{--}}{2}\ket{\emptyset}.$ (42)
The spatial mode 2 is subsequently measured in the basis of $({\hat{h}^{\dagger}}_{2}\pm{\hat{v}^{\dagger}}_{2})\ket{\emptyset}$, corresponding to the $\ket{\pm}$ of the logical qubit. If only a single $h+v$ photon is detected, the postselected output state reads
$\frac{{\hat{h}^{\dagger}}_{1}f_{++}+{\hat{v}^{\dagger}}_{1}f_{--}}{\sqrt{2}}\ket{\emptyset}.$ (43)
This is exactly the state corresponding to a graph $G^{\prime\prime}=(V^{\prime\prime},E^{\prime\prime})$ for which vertices 1 and 2 were merged with duplicate edges removed: $E^{\prime\prime}=E^{\prime}\cup\{\{1,i\}:\{1,i\}\in E\veebar\{2,i\}\in E\}$.
This is the case only if the photons in modes 1 and 2 are indistinguishable after mixing through the polarizing beamsplitter – the procedure relies on the Hong-Ou-Mandel effect. Shall this assumption not be met, the interference required for the fusion to work does not take place, and effectively the two photons originating in mode 1 and 2 are measured independently. If only one photon was observed in the new mode 2, half of the times it arrived from the mode (1,v), while the photon from mode 2 is now localized in the new mode (1,v) and the final state is ${\hat{v}^{\dagger}}_{1}f_{--}\ket{\emptyset}$. The opposite process means that the resulting state ${\hat{h}^{\dagger}}_{1}f_{++}\ket{\emptyset}$.
This effectively creates an ensemble which allows for the following interpretation: the fusion took place as if the photons were indistinguishable, but it was immediately measured in the $Z_{1}$ basis and the result of measurement is not known. This is equivalent to a simple probabilistic $Z_{1}$ flip:
$\overbrace{\Pi_{Z_{1}+}}^{\frac{1}{2}(1+Z_{1})}\rho\Pi_{Z+}+\overbrace{\Pi_{Z_{1}-}}^{\frac{1}{2}({1-Z_{1}})}\rho\Pi_{Z_{1}-}=\frac{\rho+Z_{1}\rho Z_{1}}{2}.$ (44)
The $Z$-flip ensemble is equivalent to $Z$-measurement one and is easier to work with. It is easier to observe that the noise does not propagate if the affected qubit is then fused with others, since there is always just a (potentially $Z$-flipped) graph state at the input – unless it also fails because the noise is correlated.

### v2

@@
 The discovery of cluster states – states of qubits with grid-like entanglement structure – provided a new perspective for quantum computation, better applicable for some systems [1, 2].
-The original design consisted of qubits arranged in a two-dimensional square lattice structure, where the neighboring qubits would be entangled by means of a controlled-$Z$ gate (or a specific quantum system equivalent thereof).
-Unlike the circuit-based understanding of quantum computation, the input state is constant, and the only available class of operations is the sequential measurement of individual qubits.
+The original design consisted of qubits arranged in a two-dimensional square lattice structure, where the neighboring qubits would be entangled by means of a controlled-$Z$ gates.
+Unlike the circuit-based understanding of quantum computation [3, 4], the input state is constant, and the only available class of operations is the sequential measurement of individual qubits.
 It is the act of measurement that feeds the information into the system and performs all of the needed transformations [5].
-Generalizations into different connectivity structures – the (hyper)graph states – soon followed, sharing the same core properties with the original cluster states [4, 5, 6, 7] and providing a framework for further research into the entanglement properties of multiqubit states [8, 9].
-The most basic operation in this context is no processing at all: a simple transfer or teleportation [10, 11] of one qubit state into another place.
+Generalizations into different connectivity structures – the graph and hypergraph states – soon followed, sharing the same core properties with the original cluster states [6, 7, 8, 9] and providing a framework for further research into the entanglement properties of multiqubit states [10, 11].
+Another task for which graph states are useful is the distribution of entanglement in quantum networks [12, 13, 14, 15, 16, 17].
+In the context of quantum computation, the most basic operation is no processing at all: a simple transfer or teleportation [18, 19] of one qubit state into another place.
 It is a basic building block of measurement-based quantum computation, on top of which more involved operations are built.
-In the basic scenario, this can be performed by sequential measurement of Pauli-$X$ operators on a path between the initial and end qubits (the terminal qubits in this article nomenclature) and Pauli-$Z$ on the qubits neighboring the path.
-Intuitively, the $Z$-basis measurement excises the entanglement contained in the path from the rest of the cluster state and the $X$-basis measurement along the path makes the terminal qubits correlated: final measurement of one of them fuses the two together, and the surviving qubit shares correlations with a distant part of the graph [3].
-This process can also be understood as a Bell pair generation along the path: local operations do commute, and a measurement along the path on its own does produce Bell-like correlations across the terminal qubits, which can be used directly for information transfer.
-All of this works, provided that no noise was present during the state preparation and measurement – but this is rarely the case.
-Thus, a noise-detecting or -correcting process is needed, and research in this direction has been done [12, 13, 14, 15, 16].
+In the basic scenario, teleportation can be thought of as a creation of a Bell pair, which is subsequently used to transmit the quantum state.
+The Bell pair creation step can be performed by sequential measurement of Pauli-$X$ operators on a path between the initial and end qubits and Pauli-$Z$ on the qubits neighboring the path (see Figure 1).
+Intuitively, the $Z$-basis measurement excises the entanglement contained in the path from the rest of the cluster state and the $X$-basis measurement along the path leaves the terminal (first and last) qubits in Bell-type correlations, up to local unitary rotations.
+The final measurement of one of them fuses the two together, and the surviving qubit shares correlations with a distant part of the graph ([5], see also Appendix A).
+The crucial point is that the Bell pair creation step is independent on the state to be transferred.
+The right correlational structure is prepared and subsequently used, and the preparation step can be studied alone.
+This abstraction step is helpful in various ways: the theoretical effects of the exact input state for teleportation are not relevant, this formulation allows us to utilize the works dedicated to quantum state extraction [20, 21], and similar language regarding the correlation of the relevant qubits is frequently used [22, 23], also in the research dedicated to teleportation [24].
+Figure 1: Any two vertices of a graph state associated to a connected graph can be turned into a Bell pair by sequential measurements.
+a) A grid graph describing the underlying structure of a cluster state.
+The goal is to generate a Bell pair between the qubits associated to the big red vertices.
+The chosen path is highlighted in red.
+b) We measure the qubits associated with the vertices on the path between the red vertices in $X$ basis and the qubits adjacent to the chosen path in $Z$ basis.
+c) The post-measurement state has Bell correlations between the qubits associated to the red vertices.
+This procedure was called X-protocol in [12].
 In general, entanglement is required for quantum computation, and preparation of a Bell pair or GHZ state is an operational primitive of many quantum algorithms.
 Extracting those states from larger graph states was discussed in [25, 26, 24, 13, 12].
-In this article we analyze experimentally motivated noise models and determine what can be done (to the initial noisy graph state structure and the measurement pattern) in order to mitigate the noise effects.
-This approach does not translate into full-fledged error correcting codes, but still uses the properties of the stabilizer associated with the graph state; in the ideal case, some outcomes are perfectly (anti)correlated, and any deviation indicates the presence of noise.
-As mentioned in the previous paragraph, teleportation is equivalent to a Bell pair preparation; in order to simplify the calculations, we analyze the entanglement quality of the prepared state after the internal qubits have been measured.
+The problem of extracting graph states from larger graph states is known to be NP-complete [27, 28].
+In this article we analyze experimentally motivated noise models and determine what can be done to the initial noisy graph state structure and the measurement pattern in order to mitigate the noise effects.
+This general problem of noise reduction in graph states has been analyzed from various viewpoints.
+For instance, purification and distillation protocols are discussed in [29, 30, 31].
+Quantum error correcting codes are more general, but complex.
+Unlike them, our approach utilizes only local single qubit measurements, basic preparation and verification, and is in principle adaptable to various underlying graph structures.
+However, it still uses the properties of the stabilizer associated with the graph state.
+In the ideal case, some outcomes are perfectly (anti)correlated, and any deviation indicates the presence of noise.
+As mentioned above, teleportation is equivalent to a Bell pair preparation; in order to simplify the calculations, we analyze the entanglement quality of the prepared state after the internal qubits have been measured.
 The article is structured as follows.
-In Section 2 we briefly introduce graph states and the stabilizer formalism.
-In Section 3 we explain the measurement scheme which we use to generate Bell pairs between certain nodes or for teleportation.
+In Section 2, we briefly introduce graph states and the stabilizer formalism.
+In Section 3, we explain the measurement scheme which we use to generate Bell pairs between certain nodes or for teleportation.
 We introduce noise models in Section 4 and analyze how to deal with their effects in Section 5.
-In Section 6 we show how to extend the presented methods to generate GHZ states.
+In Section 6, we show how to extend the presented methods to generate GHZ states.
 We conclude our work in Section 7.
+A portion of this article contains numerical results, the code [32] used in data generation and analysis is publicly available.
 2 Graph state preliminaries
-Figure 1: a) The square graph $G$.
-The associated state |$\ket{G}$, as any graph state, possesses an associated structure of the stabilizer $S_{G}$ composed of specific strings of Pauli operators.
-Expectation values of any Pauli string $O$ is nonzero if and only if ±$\pm O\in S_{G}$.
-$X_{1}X_{3}$ is a stabilizer operator of |$\ket{G}$ and therefore it’s expectation value is 1.
-b) Graphs corresponding to the state defined in a) being altered by noise – here, the noise may remove certain edges (dashed lines).
-Noise processes change the stabilize group (e.g.
+Figure 2: a) The square graph defines the associated graph state $\ket{\square}$ which, as any graph state, possesses an associated structure of the stabilizer $S_{\square}$ composed of specific strings of Pauli operators.
+Expectation values of any Pauli string $O$ is nonzero if and only if one of ±$\pm O$ belongs to $S_{\square}$.
+$X_{1}X_{3}$ is a stabilizer operator of $\ket{\square}$ and therefore it’s expectation value is 1.
+b) Graph states corresponding to the state defined in a) being altered by noise – here, the noise may remove certain edges (dashed lines), as in the exemplary $\ket{\sqcap}$ defined by the graph shown in the top left corner.
+Noise processes change the stabilizer group (e.g.
 by graph modification shown here), which allows for verifiable preparation.
-In this section, we present a short introduction to the class of graph states and the stabilizer formalism. Readers familiar with the topic may skip directly to the next section.
-A graph $G=(V,E)$ is defined by a set $V$ of vertices and a set of edges $E\subset V^{2}$: each $\{i,j\}=:e\in E$ denotes a connection between two vertices $i,j\in V$.
+In this section, we present a short introduction to the class of graph states and the stabilizer formalism.
+A graph $G=(V,E)$ is defined by a set $V$ of vertices and a set of edges $E\subset V^{2}$: each $\{i,j\}=:e\in E$ denotes a connection between two vertices $i,j\in V$ , where $i\neq j$.
 Graph states are multi-qubit quantum states, where the vertices and edges of the graph $G=(V,E)$ represent qubits and entangling gates, respectively.
-The state |$\ket{G}$, corresponding to a graph $G$ is defined as
+The state |$\ket{G}$, corresponding to a graph $G$, is defined as
 |$\ket{G}=\mathchoice{}{}{}{}\prod_{\{i,j\}\in E}\mathchoice{}{}{}{}\operatorname{CZ}_{i,j}\ket{+}^{\otimes\lvert V\rvert},$ (1)
-where $\operatorname{CZ}$ denotes a standard controlled-$Z$ unitary operation
+where $\operatorname{CZ}$ denotes a standard controlled-$Z$ unitary operation on qubits $i$ and $j$
 […]
 […]
-Manifestly, all of controlled-$Z$ gates are diagonal in the computation basis, and hence commute – the product in the above equation does not depend on the order of operations.
-Consider a four-vertex graph shown in Figure 1a: it has a vertex set $V=\{1,2,3,4\}$ and an edge set $E=\{\{1,2\},\{2,3\},\{3,4\},\{4,1\}\}$. The corresponding graph state is given by
-|$\ket{G}=\operatorname{CZ}_{1,2}\operatorname{CZ}_{2,3}\operatorname{CZ}_{3,4}\operatorname{CZ}_{4,1}\ket{+}^{\otimes 4}.$ (3)
+The controlled-$Z$ gates are diagonal in the computation basis, and hence commute. Therefore, the product in the above equation does not depend on the order of operations.
+Consider a four-vertex square graph shown in Figure 2a: it has a vertex set $V^{\square}=\{1,2,3,4\}$ and an edge set $E^{\square}=\{\{1,2\},\{2,3\},\{3,4\},\{4,1\}\}$.
+The corresponding graph state is given by
+|$\ket{\square}=\operatorname{CZ}_{1,2}\operatorname{CZ}_{2,3}\operatorname{CZ}_{3,4}\operatorname{CZ}_{4,1}\ket{+}^{\otimes 4}.$ (3)
 The graph states can also be described using the so-called stabilizer operators, providing a description of the state in question in terms of its correlation structure.
 Consider an initial state |$\ket{+}^{\otimes|V|}$: since it is a pure product state, there are no correlations between different sites.
-However, it is an eigenstate for each $X_{i}$ for $i\in V$.
+However, it is a $(+1)$-eigenstate for each $X_{i}$ for $i\in V$.
 Let us denote the product of controlled-$Z$ operators in Equation 1 by $U$; then
@@
 $g_{i}=X_{i}\mathchoice{}{}{}{}\prod_{j\in\operatorname{\mathcal{N}}(i)}\mathchoice{}{}{}{}Z_{j},$ (6)
-where $X_{i}$, $Y_{i}$, and $Z_{i}$ denote the Pauli matrices acting on the $i$-th qubit and $\operatorname{\mathcal{N}}(i)$ is the neighbourhood of $i$.
-As shown above, any graph state |$\ket{G}$ defined by Equation 1 is an eigenstate (to the eigenvalue of $+1$) of all $g_{i}$:
-$g_{i}\ket{G}=\ket{G}$ (7)
-While the $(+1)$-eigenspace of each of the operators $g_{i}$ is highly degenerated, taken together they fully define the state |$\ket{G}$: it is the $(+1)$-eigenstate of all $g_{i}$.
+where $X_{i}$, $Y_{i}$, and $Z_{i}$ denote the Pauli matrices acting on the $i$-th qubit and $\operatorname{\mathcal{N}}(i)$ is the neighborhood of the vertex $i$, given by
+$\operatorname{\mathcal{N}}(i)=\{j:\{i,j\}\in E\}.$ (7)
+As shown above, any graph state |$\ket{G}$ defined by Equation 1 is an eigenstate with eigenvalue $+1$ of all $g_{i}$:
+$g_{i}\ket{G}=\ket{G}.$ (8)
+While the $(+1)$-eigenspace of each of the operators $g_{i}$ is highly degenerated, taken together they fully define the state |$\ket{G}$.
+The graph state |$\ket{G}$ is the unique $(+1)$-eigenstate of all $g_{i}$.
 Operators defined by Equation 8 commute; hence, they generate an Abelian subgroup of the group of Pauli operators, called the stabilizer $S$ of |$\ket{G}$:
@@
 The elements of the group $S$, due to the compact form of the generators (Equation 6), can be described with the help of low-dimensional matrices of integers modulo $2$ [33].
-This representation is helpful in direct calculations involving the stabilizer.
+This representation is helpful in direct calculations involving the stabilizer group.
 The exemplary graph state shown in Figure 2a has stabilizer operators $g_{1}=X_{1}Z_{2}\operatorname{\mathds{1}}_{3}Z_{4}$, $g_{2}=Z_{1}X_{2}Z_{3}\operatorname{\mathds{1}}_{4}$, $g_{3}=\operatorname{\mathds{1}}_{1}Z_{2}X_{3}Z_{4}$, and $g_{4}=Z_{1}\operatorname{\mathds{1}}_{2}Z_{3}X_{4}$, as well as all products composed from these operators.
-The elements of $S$ form a complete description of all possible correlations of local measurements of graph states: if a Pauli string [^1] Here, $o_{i}$ denotes one of the Pauli operators $\{X_{i},Y_{i},Z_{i}\}$ acting on the $i$-th qubit.
-$O=o_{i_{1}}o_{i_{2}}\ldots$ (or its negative, $-O$) appears in $S$, its expectation value on |$\ket{G}$ is equal to $+1$ (respectively, $-1$), and if ±$\pm O\notin S$, $\langle O\rangle_{G}=0$.
-For example, for the graph state shown in Figure 1a, the string $O=X_{1}X_{3}=g_{1}g_{3}$ is a stabilizer operator of |$\ket{G}$.
-Using Equation 7 it can be seen that $\langle O\rangle_{G}=1$.
-The string $O^{\prime}=X_{1}Z_{3}\notin S_{G}$, hence $\langle O^{\prime}\rangle_{G}=0$.
+The elements of the stabilizer $S$ form a complete description of all possible correlations of local measurements of graph states.
+Let $O=o_{i_{1}}o_{i_{2}}\ldots$ be a Pauli string, where $o_{i}$ denotes one of the Pauli operators $\{X_{i},Y_{i},Z_{i}\}$ acting on the $i$-th qubit.
+If ±$\pm O$ appears in $S$, its expectation value on |$\ket{G}$ is equal to ±$\pm 1$.
+Otherwise, if ±$\pm O\notin S$, we have $\langle O\rangle_{G}=0$.
+For example, for the graph state shown in Figure 2a, the string $O=X_{1}X_{3}=g_{1}g_{3}$ is a stabilizer operator of $\ket{\square}$.
+Using Equation 8 it can be seen that $\langle O\rangle_{\square}=1$.
+The string $O^{\prime}=X_{1}Z_{3}\notin S_{\square}$, hence $\langle O^{\prime}\rangle_{\square}=0$.
 In this paper we are mostly using Pauli strings which contain only Pauli $X$ operators [33].
 However, the results can be generalized to other Pauli strings.
-The weighted graph states [22] are a related class of quantum states, for which controlled-phase operators are used instead of controlled-$Z$, with the phases possibly different for each edge of the graph $G$.
-If the phase associated with the edge $\{i,j\}$ is $\varphi_{i,j}$, the resulting state is defined by a modified version of Equation 1,
+Weighted graph states [34] are a related class of quantum states, for which controlled-phase operators are used instead of controlled-$Z$.
+The phases $\varphi_{i,j}$ are possibly different for each edge $\{i,j\}\in E$ of the graph $G_{\varphi}=(V,E,(\varphi_{i,j}))$.
+A modified version of Equation 1 which describes the weighted graph state is
 |${\ket{G_{\varphi}}=\mathchoice{}{}{}{}\prod_{\{i,j\}\in E}\mathchoice{}{}{}{}\operatorname{C\Phi}_{i,j}(\varphi_{i,j})\ket{+}^{\otimes\lvert V\rvert},}$ (10)
-where $\operatorname{C\Phi}(\varphi)=\operatorname{diag}(1,1,1,\exp(i\varphi))$ on the $\{i,j\}$ subsystem.S uch states arise naturally in systems with Ising-like interactions, and offer an useful representation of noisy graph state preparation using such a scheme.
+where $\operatorname{C\Phi}(\varphi)=\operatorname{diag}(1,1,1,\exp(i\varphi))$ on the $\{i,j\}$ subsystem, similarly to Equation 2.
+Such states arise naturally in systems with Ising-like interactions, and offer an useful representation of noisy graph state preparation using such a scheme.
 Unfortunately, the analogue of procedure described by Equation 5 does not produce a Pauli string, and hence a simple stabilizer description of correlations generally does not exist.
 For a standard graph state |$\ket{G}$ all phases $\varphi_{i,j}$ above are equal to $\pi$.
-This class of states naturally arises in systems with an Ising-type interaction pattern, and some physical realizations of graph states prepare the unweighted graph states in this exact way.
 Many known families of states with structured entanglement can be represented as graph states. For instance, the $N$-qubit Greenberger–Horne–Zeilinger (GHZ) state [35], defined as
-|$\ket{GHZ}=\frac{1}{\sqrt{2}}(\ket{0_{1}0_{2}\dots 0_{N}}+\ket{1_{1}1_{2}\dots 1_{N}})$ (10)
-is local unitary equivalent to the graph state of the fully connected $N$ vertex graph (or, equivalently, by the $N$ vertex star graph).
-In measurement-based quantum computing, cluster states, corresponding to grid graphs, are relevant [3], and intermediate-size cluster states have recently been realized on various platforms [24, 25].
+|$\ket{GHZ}=\frac{1}{\sqrt{2}}(\ket{0_{1}0_{2}\dots 0_{N}}+\ket{1_{1}1_{2}\dots 1_{N}}),$ (11)
+is local unitary equivalent to the graph state of the fully connected $N$ vertex graph or, equivalently, by the $N$ vertex star graph.
+In measurement-based quantum computing the cluster states, corresponding to grid graphs, are relevant [5] and are an example of a broader category of universal resource states [36].
+Intermediate-size cluster states have recently been realized on various platforms [37, 38].
 A detailed discussion of graph states, including weighted graph states, can be found in [7, 39].
 3 Correlations and measurement
-It is possible to extract an entangled pair of qubits $(i,j)$ from a graph state |$\ket{G}$ by means of local Pauli measurements, if there exists a path between $i$ and $j$ through the graph $G$ [20, 3].
-In general, there are multiple ways to perform this task – and if noise is present, the entanglement quality of the resulting two-qubit state depends on the choice.
+In this section we show how local measurement affects correlations present in graph states.
+In particular, it is possible to extract an entangled pair of qubits $(i,j)$ from a graph state |$\ket{G}$ by means of local Pauli measurements, if there exists a path between $i$ and $j$ through the graph $G$ [12, 5].
+As we later show in Section 5, the basic procedure outlined in Figure 3 is susceptible to noise effects, and the general mathematical language introduced here is used in the analysis and proposal of alternative extraction procedures.
+The post-measurement state after applying a sequence of Pauli measurements can be determined by applying graphical rules for measurements [7].
+In this section, we present an equivalent method using stabilizer operators to find the post-measurement states.
+This method is better suited for calculations.
+In general, there are multiple choices of measurement sequences.
+If noise is present, the entanglement quality of the resulting two-qubit state depends on the choice of the sequence.
 Consider a graph state |$\ket{G}$ that undergoes a local sequential measurement process involving Pauli operators.
-The state of the unmeasured qubits is completely characterized by the measurement pattern along with the outcomes; the remaining correlations stem from the stabilizer operators of |$\ket{G}$ consistent with the measurement pattern.
+The state of the unmeasured qubits is completely characterized by the measurement pattern, that is, the choice of sequential Pauli measurements, along with the outcomes.
+The remaining correlations stem from the stabilizer operators of |$\ket{G}$ consistent with the measurement pattern.
 To see the structure of these correlations, let us choose a measurement pattern on a subset of qubits $I$: for each qubit $i\in I$, a local Pauli $o_{i}\in\{X_{i},Y_{i},Z_{i}\}$ is chosen.
@@
 As mentioned previously, the expectation value of a Pauli string $O=\prod_{i\in I}o_{i}$ can be determined: if $O\in S_{G}$, it implies $\expectationvalue{O}_{G}=1$.
-The case of $-O\in S_{G}$ differs only by the sign, and if neither ±$\pm O\in S_{G}$, the expectation value is 0.
-Let us concentrate on the positive sign case (the negative sign does not change much in reasoning) and consider an actual preparation of |$\ket{G}$ and subsequent measurement according to $\{o_{i}\}_{i\in V}$: the measurement of $o_{i}$ on the $i$-th qubit has an outcome of $s_{i}$.
+The case of $-O\in S_{G}$ differs only by the sign, and if neither ±$\pm O$ belongs to $S_{G}$, the expectation value is 0.
+Without loss of generality, let us concentrate on the positive sign case and consider a preparation of |$\ket{G}$ and subsequent measurement according to $\{o_{i}\}_{i\in V}$: the measurement of $o_{i}$ on the $i$-th qubit has an outcome of $s_{i}{=\pm 1}$.
 If $\langle O\rangle_{G}=1$, the outcomes of the sequential measurement must reflect that: the signs of the measurement results must multiply to $+1$.
-Any other outcome pattern would decrease the magnitude of the experimental expectation value and is therefore incompatible with the graph state correlations.
-This reasoning applies as well if the sequential measurement is stopped at any point and resumed afterward: if $O=\prod_{i\in I}o_{i}$ is a stabilizer, and only the qubits $I^{\prime}\subset I$ were measured, the sign structure of the further measurement of $I\setminus I^{\prime}$ can be predicted.
+Any other outcome pattern would decrease the magnitude of the expectation value (defined empirically as average over random sequence of results) and is therefore incompatible with the graph state correlations.
+This reasoning applies as well if the sequential measurement is stopped at any point and resumed afterward.
+If $O=\prod_{i\in I}o_{i}$ is a stabilizer operator, and only the qubits $I^{\prime}\subset I$ were measured, the sign structure of the further measurement of $I\setminus I^{\prime}$ can be predicted.
 Thus, by the action of partial measurement, correlation in the remaining qubits is induced.
 The following Lemma captures this line of thought: the end correlations are defined by stabilizer operators consistent with the measurement scheme.
-Note that the measurement scheme does not itself have to be a stabilizer operator: only part of it has to extend to one, and this is always possible (see Lemma 3 in Appendix A).
+Note that the measurement scheme does not itself have to be a stabilizer operator: only part of it has to extend to one, and such an extension always exists (see Lemma 4 in Appendix B).
 Lemma 1.
-Let |$\ket{G}$ be a graph state determined by the graph $G=(V,E)$; we denote the stabilizer of |$\ket{G}$ by $S$.
+Let |$\ket{G}$ be a graph state determined by the graph $G=(V,E)$.
+We denote the stabilizer of |$\ket{G}$ by $S$.
 If a subset $I$ of qubits is measured in such a way that a qubit $i\in I$ is measured in the eigenbasis of $o_{i}\in\{X_{i},Y_{i},Z_{i}\}$, we encode this measurement scheme as the Pauli string $O=\prod_{i\in I}o_{i}$.
 The qubits of $I$ are measured independently and sequentially, with the measurement outcome of $o_{i}$ denoted by $s_{i}=\pm 1$.
-Let us now write a stabilizer operator $Q\in S$ of |$\ket{G}$ as a Pauli string: $Q=\pm\prod_{j\in J}q_{j}$.
+Let us now write a stabilizer operator $Q\in S$ of |$\ket{G}$ as a Pauli string with support on indices $J$: $Q=\pm\prod_{j\in J}q_{j}$, where $q_{j}$ is a Pauli operator in $\{X_{j},Y_{j},Z_{j}\}$, acting on qubit $j\in J$.
 The post-measurement state |$\ket{\psi}$ is determined by those elements of the stabilizer, which are consistent with the measurement pattern $O$ such that $Q|_{I\cap J}=O|_{I\cap J}$.
 The stabilizer operators of the post-measurement state |$\ket{\psi}$ can be computed from such operators $Q$ as
-$Q^{\prime}:=\pm\prod_{i\in I\cap J}s_{i}\prod_{j\in J\setminus I}q_{j}.$ (11)
-where the global sign is equal to the global sign of $Q$.
+$Q^{\prime}:=\pm\prod_{i\in I\cap J}s_{i}\prod_{j\in J\setminus I}q_{j},$ (12)
+where the global sign $\pm$ in front of the product is equal to the the one appearing in the definition of $Q$.
 Proof.
 Consider the stabilizer operator $Q$ defined as stated in the Lemma.
-Then assume that the measurement outcomes of each of the qubits $i\in I\cap J$ are $s_{i}$.
-Then, the projective measurement operators of these qubits can be written as $\frac{1}{2}\left(\mathbbm{1}_{i}+s_{i}q_{i}\right)$.
-[^2] For instance, the projection operator associated with the $X+$ outcome is |$\ket{X+}\bra{X+}=\frac{1}{2}\left(\mathbbm{1}+X\right)$.
+Then assume that the measurement outcomes of each of the qubits $i\in I\cap J$ is $s_{i}$.
+The projective measurement operators of these qubits can be written as $\frac{1}{2}\left(\mathbbm{1}_{i}+s_{i}q_{i}\right)$.
+For instance, the projection operator associated with the $X+$ outcome is |$\outerproduct{X+}{X+}=\frac{1}{2}\left(\mathbbm{1}+X\right)$.
 Since the projection operators commute with $q_{j}$ themselves, the following chain of equalities holds:
-|$\ket{\psi}$ $=C\prod_{i\in I\cap J}\frac{\mathbbm{1}_{i}+s_{i}q_{i}}{2}\underbrace{\ket{G}}_{\leavevmode\nobreak\ \leavevmode\nobreak\ =Q\ket{G}}$ (12)
+|$\ket{\psi}$ $=C\prod_{i\in I\cap J}\frac{\mathbbm{1}_{i}+s_{i}q_{i}}{2}\!\!\!\underbrace{\ket{G}}_{\leavevmode\nobreak\ \leavevmode\nobreak\ =Q\ket{G}}$
 $=C\prod_{i\in I\cap J}\frac{\mathbbm{1}_{i}+s_{i}q_{i}}{2}\left(\pm\prod_{j\in J}q_{j}\ket{G}\right)$
 =±$=\pm\!\!\!\prod_{j\in J\setminus I}\!\!\!q_{j}\prod_{i\in I\cap J}\!\!\!s_{i}\underbrace{\left(C\!\!\!\prod_{i\in I\cap J}\!\!\!\frac{\mathbbm{1}_{i}+s_{i}q_{i}}{2}\ket{G}\right)}_{\leavevmode\nobreak\ \leavevmode\nobreak\ =\ket{\psi}}$
-=(±$=\underbrace{\left(\pm\prod_{i\in I\cap J}s_{i}\prod_{j\in J\setminus I}q_{j}\right)}_{\leavevmode\nobreak\ \leavevmode\nobreak\ =Q^{\prime}}\ket{\psi}.$
-Here, $C$ is a normalization constant and the following equation (true for any $q$ such that $q^{2}=\mathbbm{1}$, in particular $q\in\{X,Y,Z\}$) was used:
-$\frac{\mathbbm{1}+sq}{2}q=s\frac{sq+\mathbbm{1}}{2}\quad\text{ for }s=\pm 1.$ (13)
-Figure 2: Partial local measurement of graph states induces correlations in the remaining qubits.
-Pictures is the path graph $5$ vertices: if the associated graph state is measured according to the shown measurement pattern, Bell-like correlations are induced in the two terminal qubits.
-As an example, consider the line graph of 5 qubits, as shown in Figure 2 and the Pauli string $O=X_{2}X_{3}X_{4}$ which has support on vertices in $I=\{2,3,4\}$.
-The stabilizer set of the 5 qubit line graph contains $2^{5}$ stabilizers.
-Three of them are $Q^{(a)}=X_{1}X_{3}X_{5}$, $Q^{(b)}=Z_{1}X_{2}X_{4}Z_{5}$, and $Q^{(c)}=-Y_{1}X_{2}X_{3}X_{4}Y_{5}$ which have support on qubits in $J^{(a)}=\{1,3,5\}$, $J^{(b)}=\{1,2,4,5\}$, and $J^{(c)}=\{1,2,3,4,5\}$, respectively.
-It holds that $Q^{(a)}|_{3}=O|_{3}$, $Q^{(b)}|_{2,4}=O|_{2,4}$, and $Q^{(c)}|_{2,3,4}=O|_{2,3,4}$.
-Following Lemma 1, the post-measurement state after measuring qubits 2, 3, 4 in the $X$-basis is stabilized by $s_{3}X_{1}X_{5}$, $s_{2}s_{4}Z_{1}Z_{5}$, and $-s_{2}s_{3}s_{4}Y_{1}Y_{5}$.
-This can be viewed as a graph state (Bell pair) in a nonstandard basis – the change is an effect of the measurement performed on the qubits $\{2,3,4\}$.
+=(±$=\underbrace{\left(\pm\!\!\!\prod_{i\in I\cap J}s_{i}\prod_{j\in J\setminus I}\!\!\!q_{j}\right)}_{\leavevmode\nobreak\ \leavevmode\nobreak\ =Q^{\prime}}\ket{\psi}.$ (13)
+Here, $C$ is a normalization constant and the following property
+$\frac{\mathbbm{1}+sq}{2}q=s\frac{sq+\mathbbm{1}}{2}\quad\text{ for }s=\pm 1$ (14)
+was used, valid for any $q$ such that $q^{2}=\mathbbm{1}$ and in particular for $q\in\{X,Y,Z\}$. ∎
+Figure 3: Partial local measurements on a graph states induces correlations in the remaining qubits.
+a) If the 4 vertex path graph state is measured according to the shown measurement pattern, that is, measuring the inner vertices in the $X$-basis, we get a connected two-vertex graph.
+b) The graph here is not connected: the edge $\{2,3\}$ is absent, possibly due to scenario outlined in Section 4.1.
+If the corresponding graph state is measured according to the same measurement pattern, we get a disconnected two-vertex graph.
+Measurements can never induce entanglement between parties that were separable before.
+As an example, consider the path graph of 4 qubits, as shown in Figure 3a and the Pauli string $O=X_{2}X_{3}$ which has support on the vertices in $I=\{2,3\}$.
+The stabilizer of the four-qubit path graph contains $2^{4}$ stabilizer elements.
+Three of them are $Q^{(a)}=X_{1}X_{3}Z_{4}$, $Q^{(b)}=Z_{1}X_{2}X_{4}$, and $Q^{(c)}=-Y_{1}X_{2}X_{3}Y_{4}$ which have support on qubits in $J^{(a)}=\{1,3,4\}$, $J^{(b)}=\{1,2,4\}$, and $J^{(c)}=\{1,2,3,4\}$, respectively.
+We have $Q^{(a)}|_{3}=O|_{3}$, $Q^{(b)}|_{2}=O|_{2}$, and $Q^{(c)}|_{2,3}=O|_{2,3}$.
+[^1] The notation $O|_{S}$ denotes restriction of a product of operators to a qubit subset $S$.
+Following Lemma 1, the post-measurement state after measuring qubits 2 and 3 in the $X$-basis is stabilized by $s_{3}X_{1}Z_{4}$, $s_{2}Z_{1}X_{4}$, and $-s_{2}s_{3}Y_{1}Y_{4}$.
+For $s_{2},s_{3}=+1$, these are stabilizer operators of the two-vertex graph state with one edge between them, as shown in the lower part of Figure 3a.
+For other measurement outcomes $s_{2},s_{3}$, we get stabilizer operators which are equivalent up to local unitary operations.
+Recall that this graph state is local unitary equivalent to the Bell states.
+Another example of a four-vertex disconnected graph is shown in Figure 3b.
+This graph state is also stabilized by the $2^{4}$ operators.
+The three of them that coincide with the measurement pattern $O=X_{2}X_{3}$ are $Q^{(a)}=Z_{1}X_{2}$, $Q^{(b)}=X_{3}Z_{4}$, and $Q^{(c)}=Z_{1}X_{2}X_{3}Z_{4}$ that have support in $J^{(a)}=\{1,2\}$, $J^{(b)}=\{3,4\}$, and $J^{(c)}=\{1,2,3,4\}$, respectively.
+We have $Q^{(a)}|_{2}=O|_{2}$, $Q^{(b)}|_{3}=O|_{3}$, and $Q^{(c)}|_{2,3}=O|_{2,3}$.
+Following Lemma 1, the post-measurement state after measuring qubits 2 and 3 in the $X$-basis is stabilized by $s_{2}Z_{1}\operatorname{\mathds{1}}_{4}$, $s_{3}\operatorname{\mathds{1}}_{1}Z_{4}$, and $s_{2}s_{3}Z_{1}Z_{4}$.
+No correlations are present, since both remaining qubits are stabilized by independent local operators, and the resulting state is a product one.
+Note that the stabilizer operators of post-measurement states are not always graph state stabilizer operators, but are always locally unitarily equivalent to them [7].
+The example in Figure 3a is chosen so that the post-measurement state does not require this local unitary transformations.
+Choosing a path graph of odd length results in different Bell state stabilizers.
+In the example shown in Figure 3b, we get compositions of Pauli $Z$-operators as post-measurement stabilizers.
+Applying the local unitary $H_{1}H_{4}$ (where $H$ is the Hadamard gate) to them leads to the actual graph state stabilizer operators.
 The remaining qubits are fully defined by the correlations developed during the measurement process.
-Thus, it is possible to find measurement schemes for which the output state has certain properties: e.g., if the goal is to extract a Bell pair from a larger graph state |$\ket{G}$, a measurement scheme $\prod_{i\in I}o_{i}$ must be found such that there exist stabilizers $Q$ of |$\ket{G}$ such that $Q_{I\cup J}=O_{I\cup J}$ and ensuring Bell-like correlations between the terminal vertices.
-For example in the 5 qubit line graph and the above chosen measurements, the Bell state stabilizers are extracted; more elaborate measurement patterns have been studied for use in quantum networks [20, 19, 27].
-Further examples for graph families analyzed later in this article can be found in Appendix A.
+Thus, it is possible to find measurement schemes for which the output state has certain properties.
+For example, if the goal is to extract a Bell pair from a larger graph state |$\ket{G}$, a measurement scheme $\prod_{i\in I}o_{i}$ must be found such that there exist stabilizers $Q$ of |$\ket{G}$ such that $Q_{I\cap J}=O_{I\cap J}$ and ensuring Bell-like correlations between the terminal vertices.
+For example in the 4-qubit path graph of Figure 3a and the above chosen measurements, the Bell state stabilizers are extracted.
+More elaborate measurement patterns have been studied for use in quantum networks [12, 13, 40].
+Further examples for graph families analyzed later in this article can be found in Appendix B.
 As a direct result of the Lemma 1 we get the following observation:
 Corollary 2.
-If a stabilizer operator $Q=\pm\prod_{j\in J}q_{j}$ is fully embedded in the measurement scheme $O=\prod_{i\in I}o_{i}$ such that $J\subset I$ and $Q=\pm O|_{J}$, the only observable sign structures within $J$ are those such that $\prod_{j\in J}s_{j}=\pm 1$, where the sign is determined by the sign in $Q$.
+If a stabilizer operator $Q=\pm\prod_{j\in J}q_{j}$ is fully embedded in the measurement scheme $O=\prod_{i\in I}o_{i}$ such that $J\subset I$ and $Q=\pm O|_{J}$, the only observable sign structures within $J$ are those such that $\prod_{j\in J}s_{j}=\pm 1$, where the sign is determined by the sign $\pm$ in front of the product defining $Q$.
 The proof follows by considering only $J\subset I$ in Lemma 1.
-As mentioned in the introduction, the creation of a Bell pair is equivalent to the task of information transfer (teleportation) across the graph.
-This is so because if the Bell pair creation in the graph $G$ is viewed in the context of a larger graph $G^{\prime}$, this process (followed by measurement of one of the terminal qubits) performs qubit fusion: the unmeasured terminal qubit acts as the two combined.
-Since all local measurements with nonoverlapping supports do commute, this can be done before any other measurement or afterwards: in the latter case, the information created in one of the qubits is merged with the second one.
-This is captured by the following Lemma, represented pictorially by the Figure 3.
-Figure 3: If a measurement scheme generates a Bell pair across the vertices $\{{\operatorname{in}},{\operatorname{out}}\}$ in the state corresponding to a graph $G$, in a larger $G^{\prime}$ it can be used to transfer information across the graph.
-Extending the measurement by a properly chosen operator $A_{{\operatorname{in}}}$ leads to the fusion of the two terminal vertices.
-Lemma 2.
-Consider a graph $G=\{V,E\}$, and choose the two terminal vertices $\{{\operatorname{in}},{\operatorname{out}}\}$ such that $V=\{{\operatorname{in}},{\operatorname{out}}\}\cup I$.
-Suppose that the corresponding graph state |$\ket{G}$ undergoes a sequential measurement of $I$ producing a Bell pair across $\{{\operatorname{in}},{\operatorname{out}}\}$ in the meaning of Lemma 1.
-With this determined, let us further assume that $G^{\prime}=(V^{\prime},E^{\prime})$ embeds $G$ in such a way that $\{{\operatorname{in}},{\operatorname{out}}\}$ is a separating set: the only connections between $G$ and the rest of $G^{\prime}$ are at these two vertices.
-Let us denote the vertices in $V^{\prime}\setminus V$ to which both qubits are connected by $\mathcal{N}^{\prime}({\operatorname{in}})$ and $\mathcal{N}^{\prime}({\operatorname{out}})$, respectively, and assume that the two sets are disjoint.
-Within $G^{\prime}$, the aforementioned Bell pair-generating measurement followed by a product of single qubit unitary operations on the qubits $\{{\operatorname{in}},{\operatorname{out}}\}$ and the measurement $X_{\operatorname{in}}$ effectively performs qubit fusion of the terminal qubits.
-The new state is a graph state where the qubits $V^{\prime}\setminus V$ are intact, and the new neighborhood of the remaining qubit is $\mathcal{N}_{{\operatorname{out}}}\cup\mathcal{N}_{{\operatorname{in}}}$
-Proof.
-If within $G$ a Bell pair is produced across $\{{\operatorname{in}},{\operatorname{out}}\}$ as a result of the measurement pattern $O=\prod_{i\in L}o_{i}$ (where $L=V\setminus\{{\operatorname{in}},{\operatorname{out}}\}$), this means that there exist three $O$-consistent stabilizers involving the terminal vertices in the sense of Lemma 1.
-Combinatorical considerations show that one of those must be of a form such that the operators at the terminal vertices are $A_{\operatorname{in}},B_{\operatorname{out}}\in\{X,Y\}$ (but not $Z$), chosen independently.
-Thus, there exists an $G$-stabilizer operator of the form ±$\pm A_{\operatorname{in}}B_{\operatorname{out}}\prod_{i\in I}o_{i}=g_{\operatorname{in}}g_{\operatorname{out}}\prod_{i\in I^{\prime}}g_{i}$, with $I,I^{\prime}\subset L$.
-Note that the definition of $g_{i}$ depends on the graph in question (Equation 6): if by $g^{\prime}_{i}$ we denote the generators corresponding to the entire graph $G^{\prime}$, the result is also a stabilizer, differing only by the $Z$ operators in the set $\mathcal{N}^{\prime}({\operatorname{in}})\cup\mathcal{N}^{\prime}({\operatorname{out}})$.
-After the $A_{\operatorname{in}}$-basis measurement, we get a new stabilizer provided by Lemma 1:
-(±$\left(\pm s_{\operatorname{in}}\prod_{i\in L}s_{i}\right)B_{\operatorname{out}}\mathchoice{}{}{}{}\prod_{j\in\mathcal{N}^{\prime}_{\operatorname{in}}\cup\mathcal{N}^{\prime}_{\operatorname{out}}}\mathchoice{}{}{}{}Z_{j},$ (14)
-which can be taken to be a new stabilizer generator associated with the vertex ${\operatorname{out}}$, and brought to the canonical form of $X_{\operatorname{out}}\prod_{j\in\mathcal{N}^{\prime}_{\operatorname{in}}\cup\mathcal{N}^{\prime}_{\operatorname{out}}}Z_{j}$ by local unitary basis change of the ${\operatorname{out}}$-qubit.
-The properties of other stabilizers involving the qubits $\{{\operatorname{in}},{\operatorname{out}}\}$ can be proven similarly.
-Thus, after the measurement within the Bell-generating part $L\cup\{{\operatorname{in}}\}$, the ${\operatorname{out}}$-qubit behaves exactly like it was connected to the neighborhood $\mathcal{N}^{\prime}_{\operatorname{in}}$ as well: the two vertices are fused and the information between them is transferred.
-∎
 4 Imperfect graph states
-The reasoning presented above assumes that perfect $\operatorname{CZ}$ gates can be performed.
-However, this is not the case in experimental setups: some kind of noise is always present.
-In this section, we analyze three classes of physically motivated noise models; they all stem from different methods of graph state preparation.
-The origins and effects are sketched; for the detailed derivation, please refer to Appendix B.
-Regardless of origin, noise processes often have a stabilizer-consistent description: vaguely speaking, they produce a mixture of stabilizer states, or the noise effect can be otherwise understood using the stabilizer-like correlations.
-In later sections, we show how to use this observation to mitigate the effects of noise.
-4.1 Uncorrelated edge noise
-Implementations like superconducting qubit quantum computers [28, 29], or ion traps [30, 31] prepare graph states by letting an initial product state evolve via an engineered Ising-like interaction Hamiltonian:
-$H=\sum_{i\neq j}\gamma_{i,j}\frac{(Z_{i}+\operatorname{\mathds{1}})\otimes(Z_{j}+\operatorname{\mathds{1}})}{4}.$ (15)
+In this section, we analyze three classes of physically motivated noise models.
+They all stem from different methods of graph state preparation.
+The origins and effects are only sketched here due to complexity, please refer to Appendix C for the detailed derivation.
+As outlined in the previous section, for perfect graph states Bell pairs can be extracted, provided the vertices of interest are connected by a path.
+Once a path is found, vertices adjacent to it are measured in the $Z$-basis and then the path vertices in the $X$-basis.
+In Section 5 we show that in the presence of noise, the performance of a simple path graph is not optimal, and choice of other graphs can enhance entanglement quality of the remaining qubits.
+As a minimal example, consider the graph shown in Figure 3b.
+It may arise during preparation of path graph (Figure 3a) via an imperfect $\operatorname{CZ}$ gate, and in such case the Bell pair can not be generated: the two graph parts are uncorrelated, and local measurements can not produce entanglement from a product state.
+In a path graph, even a single missing edge leads to this scenario.
+Regardless of origin, noise processes often have a stabilizer-consistent description.
+They produce a mixture of stabilizer states, or the noise effect can be otherwise understood using the stabilizer-like correlations.
+In Section 5, we show how to use this observation to mitigate the effects of noise.
+Implementations like superconducting qubit quantum computers [41, 42], or ion traps [43, 44] prepare graph states by letting an initial product state evolve via an engineered Ising-like interaction Hamiltonian
+$H=\sum_{i\neq j}\gamma_{i,j}\frac{(Z_{i}+\operatorname{\mathds{1}})\otimes(Z_{j}+\operatorname{\mathds{1}})}{4},$ (15)
+where $\gamma_{i,j}$ is the interaction strength between qubits $i$ and $j$.
 Evolution according to such an interaction pattern effectively implements a sequence of controlled-phase operators $\operatorname{C\Phi}_{i,j}=\operatorname{diag}(1,1,1,\exp(-i\varphi_{i,j}))$.
-The structure of phases $\varphi_{i,j}=\gamma_{i,j}t$ is determined by the interaction strengths $\gamma_{i,j}$ and the evolution time $t$: inaccuracies in controlling either of them lead to the generation of a weighted graph state [22].
+The structure of phases $\varphi_{i,j}=\gamma_{i,j}t$ is determined by the interaction strengths $\gamma_{i,j}$ and the evolution time $t$.
+Inaccuracies in controlling either of them lead to the generation of a weighted graph state [34].
 However, under general assumptions, the effective quantum channel implemented this way has a simple decomposition into local (one or two edges) operations, on top of the product of a sequence of $\operatorname{CZ}$ operations preparing the desired state.
@@
 We discuss the two extremal cases of correlations of phase noises $\varepsilon_{i,j}$.
-The first, uncorrelated phase noise, assumes that the noises at different edges $(i,j)$ are completely independent.
+The first, uncorrelated phase noise, assumes that the noises at different edges $\{i,j\}$ are completely independent.
 The other, correlated phase noise, assumes the contrary: each phase noise factor is equal to any other in a single experimental run, $\varepsilon_{i,j}=\varepsilon$.
 The noise encountered in experiments likely has yet a different structure with partial correlations between the edges; still, the two extremal processes may help in modeling the real-life scenario and strategies developed to mitigate them should apply in the more general cases.
-Phases uncorrelated across the edges effectively lead to a probabilistic $\operatorname{CZ}$ (see the Appendix B) gate being implemented, if only each $\varepsilon_{i,j}$ is symmetrically distributed around 00 (so that $\varphi=\pi+\varepsilon$ is centered around $\pi$).
+4.1 Uncorrelated edge noise
+The effect of phase noise is especially simple if the phases $\varphi$ are distributed symmetrically and uncorrelated around $\pi$.
+The resulting channel is described by a probabilistic application of $\operatorname{CZ}$: with probability $(1-p)$, an edge is created via $\operatorname{CZ}$, and with probability $p$, no operation is performed.
 Each of controlled-$Z$ unitary operators in Equation 1 is replaced with the quantum channel
 $\operatorname{\mathcal{CZ}_{p}}[\rho]=(1-p)\operatorname{\mathcal{CZ}}[\rho]+p\rho,$ (16)
-where $p$ depends on the distribution of $\varepsilon$ and $\operatorname{\mathcal{CZ}}[\rho]=\operatorname{CZ}\rho\operatorname{CZ}^{\dagger}$.
-Thus effectively the effect of noise is the generation of randomized graph states [32]: ensemble of graph states built atop the original graph $G$ by removing the edge $(i,j)\in E$ with probability $p_{i,j}$; related graph ensembles have been introduced in [33] in a classical context.
-For the constant probability, the end state can be written as
-$\rho=\sum_{E^{\prime}\subset E}(1-p)^{|E^{\prime}|}p^{|E\setminus E^{\prime}|}\ket{G^{\prime}}\bra{G^{\prime}},$ (17)
+[…]
+$\rho=\sum_{E^{\prime}\subset E}(1-p)^{|E^{\prime}|}p^{|E\setminus E^{\prime}|}\outerproduct{G^{\prime}}{G^{\prime}},$ (17)
 where $G^{\prime}=(V,E^{\prime})$ is a graph with edge set $E^{\prime}$ being a subset of the edge set $E$ of the original graph $G$.
-Note that it is an effective description: in each experimental run, the prepared state is a pure weighted graph state [13]; the density operator in Equation 17 is a state of knowledge about the system, averaged over different noise realizations.
+Note that it is an effective description: in each experimental run, the prepared state is a pure weighted graph state [21].
+The density operator in Equation 17 is a state of knowledge about the system, averaged over different noise realizations.
 4.2 Correlated edge noise
-If the weights in the Hamiltonian of Equation 15 are perfect ($\gamma_{i,j}=1$) but the interaction time is not perfectly controlled, within one experimental run, all the resulting phases $\varphi_{i,j}$ are equal.
-This is correlated phase noise: it has an effect similar, but more involved, to the uncorrelated case.
+If the weights in the Hamiltonian of Equation 15 are all equal ($\gamma_{i,j}=1$) but the interaction time is not perfectly controlled, within one experimental run, all the resulting phases $\varphi_{i,j}$ are equal.
+This is the correlated phase noise: it has an effect similar, but more involved, to the uncorrelated case.
 Here, we also assume that $\varphi$ is symmetrically distributed around $\pi$.
 Let us assume that in each experimental run $\varphi_{i,j}=\varphi=\pi+\varepsilon$, and $\varepsilon$ is distributed with the normal distribution of the standard deviation $\sigma$.
-Then, with $p$ depending on the phase distribution (see Appendix B for details), the lowest order of approximation of the state (essentially, keeping only terms of order $\varepsilon^{2}$ in the end state and subsequent averaging) can be described with the help of square roots of the unitary operator $\operatorname{CZ}$:
-$\rho=$ $(1-p\lvert E\rvert)\ket{G}\bra{G}+p\sum_{e\in E}\outerproduct{G-e}{G-e}$ | (18)
-$+p\!\!\!\!\sum_{\{e,e^{\prime}\}\subset E\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \phantom{x}}^{e\neq e^{\prime}}\!\!\!\!\!\!\!\!\!\!\!\!\!\sum_{\phantom{x}\leavevmode\nobreak\ \leavevmode\nobreak\ \pm_{1},\pm_{2}}\!\!\!\!\!\!\!\pm_{1}\!\pm_{2}\!{\operatorname{\mathcal{CS}}_{e}^{\pm_{1}}}\!\!{\operatorname{\mathcal{CS}}_{e^{\prime}}^{\pm_{2}}}[\ket{G}\bra{G}].$
-Note that this decomposition is not convex and cannot be interpreted as an ensemble: the minus signs in front of the decomposition terms prevent this.
-Signs appear as a result of discrete decomposition of controlled-phase and subsequent averaging; see Appendix B for details.
-In this decomposition, $G-e\coloneqq(V,E\setminus\{e\})$ is the graph with edge removed, and ${\operatorname{\mathcal{CS}}^{\pm}}$ is the application of one of the two unitary square roots of $\operatorname{CZ}$:
-${\operatorname{\mathcal{CS}}^{\pm}}[\rho]=\sqrt{\operatorname{CZ}^{\pm}}\rho\sqrt{\operatorname{CZ}^{\pm}}^{\dagger},$ (19)
-where
-$\sqrt{\operatorname{CZ}^{\pm}}\coloneqq\frac{1}{2}\left[\left(1+e^{\pm i\frac{\pi}{2}}\right)\operatorname{\mathds{1}}+\left(1-e^{\pm i\frac{\pi}{2}}\right)\operatorname{CZ}\right].$ (20)
+The resulting state can be described by a parameter $p$ depending on the distribution of $\varepsilon$ (Equation 34).
+By keeping only the lowest nontrivial order of noise effects and subsequent averaging, the end state can be described with the help of square roots of the unitary operator $\operatorname{CZ}$:
+[…]
+$+\frac{p}{2}\sum_{\begin{subarray}{c}e,e^{\prime}\in E\\ e\neq e^{\prime}\end{subarray}}\sum_{s,s^{\prime}\in\{\pm 1\}}ss^{\prime}\operatorname{\mathcal{CS}}_{e}^{s}\operatorname{\mathcal{CS}}_{e^{\prime}}^{s^{\prime}}[\outerproduct{G}{G}],$
+where the summation runs over graphs $G^{\prime}\subseteq G$ with single edges removed from $G$, and ${\operatorname{\mathcal{CS}}^{\pm}}[\rho]$ is the application of one of the two unitary square roots of $\operatorname{CZ}$.
+See Appendix C for details.
 4.3 Local $Z$ flip noise
-In linear optic experiments, where Bell pairs are generated and subsequently fused to get a graph state, the entangling operation may fail as a result of partial photon distinguishability (by imperfect frequency of spatial mode overlap) [34, 35, 36] – for a derivation see the Appendix B.
+In linear optic experiments, where Bell pairs are generated and subsequently fused to get a graph state, the entangling operation may fail as a result of partial photon distinguishability This can be caused by imperfect frequency or spatial mode overlap [47, 48, 49].
+See LABEL:{app:noise} for a derivation.
 The effect of a noisy fusion gate described this way can be modeled as a perfect fusion followed by a probabilistic application of the $Z$ unitary to the surviving optical qubit $i$:
@@
 4.4 Quantification of the noise effects
-In the following chapters, we determine the entanglement quality of a small (two- or three-qubit) subsystem left after a measurement procedure performed on a graph state: this is the Bell pair or GHZ state extraction, since ideally the resulting state would represent correlations found in these states.
-We have chosen fidelity with respect to the ideal state as a measure of the entanglement quality.
-In the absence of noise, the initial state |$\ket{G}$ is pure, and measurement of an $i$-th qubit in the basis of an operator $o_{i}\in\{X_{i},Y_{i},Z_{i}\}$ leads to observation of an outcome ±$\pm 1$.
-On the algebraic level, this corresponds to application of an operator $\frac{1}{2}\left(\mathbbm{1}\pm o_{i}\right)$ to the pre-measurement state, and after the entire procedure, the post-measurement state |$\ket{\psi}$ is still pure.
-This allows a simple determination of the fidelity with respect to the noisy state $\rho$: if the post-measurement state is $\sigma$, the following holds.
+In the following sections, we determine the entanglement quality of two- and three-qubit Bell and GHZ graph states left after a measurement procedure performed on a graph state.
+To quantify the entanglement quality of the post-measurment state, we use the fidelity with respect to the ideal state.
+If the post-measurement state is $\sigma$, the following holds:
 (|$\mathcal{F}(\ket{\psi}\!\bra{\psi},\sigma)=\langle\psi|\sigma|\psi\rangle.$ (20)
 Note that the ideal post-measurement state depends on the observed outcomes during the experiment: they do affect the signs in the final stabilizer structure (see Lemma 1).
-Therefore, the ideal state |$\ket{\psi}$ must be determined from the signs of the outcomes.
-This is only possible if the structure is consistent with the stabilizers of the ideal state; this is not an issue, since we explicitly postselect on the right set of outcomes in the following sections as part of the error mitigation strategy.
-Finally, the average fidelity $\langle\mathcal{F}\rangle$ is determined for all possible measurement outcomes consistent with the stabilizer structure; this is effectively approximated by taking $N$ samples from the defining ensembles for a specific noise model with a Monte Carlo algorithm and thus
-$\langle\mathcal{F}\rangle=\frac{1}{N}\!\!\!\sum_{\text{outcomes}}\!\!\!\mathcal{F}(\text{outcome}).$ (23)
-Since we are mostly interested in the behavior in the low noise limit, we define fidelity susceptibility $\alpha$ as the rate of change of mean fidelity $\langle\mathcal{F}\rangle(p)$: the ensemble of extracted and postselected states depends on the noise, and the susceptibility captures the effects of small but not negligible noise.
+Therefore, the ideal state |$\ket{\psi}$ must be determined from the signs of the outcomes, provided they are consistent with the stabilizer structure.
+The different outcomes may appear with varying probabilities, and the expectation value of associated fidelities (Equation 20) is the average fidelity $\langle\mathcal{F}\rangle$.
+This is effectively approximated by taking $N$ samples from the defining ensembles for a specific noise model with a Monte Carlo algorithm and thus
+$\langle\mathcal{F}\rangle\approx\frac{1}{N}\!\!\!\sum_{\text{outcomes}}\!\!\!\mathcal{F}(\text{outcome}).$ (21)
+Since we are mostly interested in the behavior in the low noise limit, we define fidelity susceptibility $\alpha$ as the rate of change of mean fidelity $\langle\mathcal{F}\rangle(p)$: the ensemble of extracted and postselected states depends on the noise, and the susceptibility captures the effects of small but not negligible noise:
 […]
@@
 For any other amount, the fidelity may only decrease, and hence the susceptibility is positive.
-Its numerical magnitude determines how fragile the extraction procedure is to the effects of noise: a robust one has a small (or, in some cases, even zero) susceptibility.
+Its numerical magnitude determines how fragile the extraction procedure is to the effects of noise: a robust one has a small susceptibility.
+Practically, the fidelity susceptibility is estimated here by a discrete derivative approximation.
+The fidelity $\langle\mathcal{F}\rangle_{*}$ is determined by choosing a small value of noise $p_{*}$ and computing the fidelity for it.
+Finally, the fidelity susceptibility is approximated as $\alpha\approx\frac{1-\langle\mathcal{F}\rangle_{*}}{p_{*}}$.
+In our computations, we chose $p_{*}\approx 10^{-2}$.
+The noise value $p_{*}$ were chosen such that it is empirically low enough for reasonable approximation and high enough such that numerical noise is not relevant.
 5 Extraction from noisy ensembles
 In this section we apply the methods developed in previous parts to show the effects of the noise and find measurement patterns that minimize them.
-Instead of focusing on teleportation directly, we analyze the (equivalent) task of extracting a high-quality Bell pair from a larger noisy state, close to a graph state |$\ket{G}$.
+As mentioned previously, we analyze task related to teleportation: extraction of a high quality Bell pair from a larger noisy state, close to a graph state |$\ket{G}$.
+Various graph families are introduced together with appropriate measurement patterns, and the effect of different types of noise is analyzed for all of them.
+The main results are shown in Figures 6 and 7.
+Monte Carlo sampling was used to calculate the numerical results presented in this section.
+We mostly concentrate on the scenario where the entangling gates are almost perfect, and in such case the fidelity susceptibility (Equation 22) quantifies the noise effects.
+Practically, the average fidelity (Equation 21) is determined by simulating the random outcomes of quantum measurement (with postselection where appropriate), and so is the fidelity susceptibility $\alpha$.
+For the estimates of $\alpha$ only a limited number of quantum states forming the ensemble (discussed in Section 4) is relevant, and for calculation only this subset is taken into account, reducing the uncertainty inherent to Monte Carlo procedure.
+The code used for calculation is publicly available [32], the numerical details can be found there.
 5.1 Noise-correcting structures and robust families of states
-As introduced above, noise processes lead to an ensemble of different states $\{(p_{i},\ket{\psi_{i}}\}$ being prepared, where one of the ensemble components is the ideal state |$\ket{G}$.
+As introduced in Section 4, noise processes lead to an ensemble of different states $\{(p_{i},\ket{\psi_{i}})\}$ being prepared, where one of the ensemble components is the ideal state |$\ket{G}$.
 This ensemble can be interpreted as a probabilistic preparation of an unknown state |$\ket{\psi_{i}}$ with probability $p_{i}$.
-The same measurement procedure as previously might now yield results inconsistent with |$\ket{G}$, and observations of those are an indication of state other than |$\ket{G}$ being prepared.
-As an example, consider the graph $G$ to be a $4$-cycle of vertices $\{1,2,3,4\}$, as shown in Figure 1a.
-The operator $O=X_{1}X_{3}$ belongs to a stabilizer of |$\ket{G}$, thus the only sign structures that can be observed in the sequential measurement of qubits $1$ and $3$ are $+1,+1$ and $-1,-1$.
-Consider now all edge subgraphs $\{G^{\prime}\}$ of $G$ and a physical process in which an ensemble of graph states $\{p_{G^{\prime}},\ket{G^{\prime}}\}$ is prepared, some of them are shown in Figure 1b.
-The operator $O$ (and $-O$) is not a part of the stabilizer of any proper subgraph with at least one edge; thus, sign structures of $+1,-1$ or $-1,+1$ are possible: observing those in the same measurement procedure implies that the pre-measurement state was not |$\ket{G}$.
+The measurement procedure might now yield results inconsistent with |$\ket{G}$, and observations of those are an indication of state other than |$\ket{G}$ being prepared – see Corollary 2.
+As an example, consider the square graph – a $4$-cycle of vertices $\{1,2,3,4\}$, as shown in Figure 2a.
+The operator $O=X_{1}X_{3}$ is an element of the stabilizer of $\ket{\square}$, thus the only sign structures that can be observed in the sequential measurement of qubits $1$ and $3$ are $+1,+1$ and $-1,-1$.
+Consider now all edge subgraphs $\{G^{\prime}\}$ of the square graph and a physical process in which an ensemble of graph states $\{p_{G^{\prime}},\ket{G^{\prime}}\}$ is prepared, some of them are shown in Figure 2b.
+The operator $O$ (and $-O$) is not a part of the stabilizer of any proper subgraph with at least one edge; thus, sign structures of $+1,-1$ or $-1,+1$ are possible.
+Observing those in the same measurement procedure implies that the pre-measurement state was not $\ket{\square}$.
 Figure 4: Considered families of graphs for the task of Bell pair extraction, parameterized by the length of the internal part $n$.
-The terminal qubits, designated for creation of a Bell pair, are denotes with filled nodes.
+The terminal qubits, designated for creation of a Bell pair, are denoted with filled nodes.
 From the top: path graph, twisted pair, crazy graph.
-The postselection on the right measurement outcomes consistent thus performs a probabilistic check whether |$\ket{G}$ was truly prepared.
-However, the existence of such parity-checking structures depends on the graph and the task for which the associated quantum state is used.
-A Bell pair can be prepared with the help of a path graph state and local $X$ measurements, but such a system does not support embedded stabilizer operators.
-Other graphs and measurement patterns do better in this regard, and here we present them.
+Therefore, the postselection on the right measurement outcomes performs a probabilistic check whether |$\ket{G}$ was truly prepared.
+However, the existence of such parity-checking structures depend on the graph and the task for which the associated quantum state is used.
+A Bell pair can be prepared with the help of a path graph state and local $X$ measurements (Figure 3a), but such a system does not support embedded stabilizer operators.
+Other graphs and measurement patterns do better in this regard.
+Here we present some alternatives.
 The families of graph states analyzed are parameterized by the length between the terminal qubits across which a Bell pair is to be prepared (see Figure 4).
 Apart from a simple path graph, included as a benchmark for comparison, we include two other families described below – both of them have their strengths.
-The twisted pair graph [37] is built from layers, where each qubit in the layer $k$ is connected with every qubit from the layers $k-1$ and $k+1)$.
+The twisted pair graph [50], shown in Figure 4b, is built from layers, where each qubit in the layer $k$ is connected with every qubit from the layers $k-1$ and $k+1$.
 The number of qubits in each layer alternates between $1$ and $2$, which enables the existence of a restricted set of stabilizers: for each 2-qubit layer consisting of the qubits $\{i,j\}$, the operators $X_{i}X_{j}$ stabilize |$\ket{G}$.
 Additionally, it is very robust to certain types of noise, because of its relatively simple geometry.
-The crazy graph (studied in [38, 14] for its noise robustness) also consists of layers of $2$ qubits each, with a similar full connectivity structure between the adjacent layers.
-This ensures a robust and simple to analyze structure of the embedded stabilizer operators (see Appendix A and Figure 5).
+The crazy graph similarly consists of layers of $2$ qubits each, with full connectivity between the adjacent layers.
+It is known for its noise robustness [51, 24], which stems from the structure of its stabilizer operators.
+See Figure 5 and Appendix B for more details.
 Figure 5: Crazy graph of length 3 together with an all-$X$ measurement pattern possess embedded error-checking and correction structures.
@@
 Bottom: even if any single internal edge (dotted line between layers 1 and 2) is missing, the terminal qubits correlations are not affected: the operators shown are stabilizers, regardless of whether the edge is present or not.
-Both presented structures generalize to the extraction of other types of states, in which context they serve as a direct replacement of path graphs.
-We mention the usage in a task of 3-qubit GHZ state extraction.
-5.2 Bell pair quality scaling for different type of noise
-Each of the graph states associated with the aforementioned three families (path graph, twist graph, and crazy graph) can be used for extraction of a Bell pair across the terminal qubits.
-Only local $X$ measurements are used for all three, and the fact that in the ideal case a maximally entangled state is produced is captured by the stabilizer operators consistent with this measurement pattern.
-As a simple example, consider the 5 qubit line graph as discussed in Section 3.
-Measurements in the $X$ basis on all but the terminal qubits lead to a two-qubit state which can be shown to have Bell-like correlations.
-Similar structures can be found for all the three graph families mentioned: sets of stabilizer operators ensuring correlations of the terminal qubits, having only $X$ operators in the internal section (and thus compatible with the measurement scheme).
-The effect of noise destroys the perfect correlations: even if for every ensemble element the end state is maximally entangled, the mixture of such states might have less (or none) entanglement.
-Postselection on the measurement outcomes consistent with the ideal stabilizer structure helps twofold: one of its effects is probabilistic error detection.
-The other, appearing in the case of the crazy graph ensemble, is more subtle: after postselection, the end state may be the same for a large portion of the ensemble elements (see Appendix A).
-The results, presented in Figure 6, show that this approach (noise-detecting structures combined with postselection) works for different types of noise.
-The fidelity susceptibility (quantifying the initial fidelity falloff in the low noise limit; see Equation 24) is reduced for the twisted pair and crazy structures.
+Each of the graph states associated with the aforementioned three families (path graph, twist graph, and crazy graph) can be used for extraction of a Bell pair across the terminal qubits (shaded vertices in Figure 4).
+Only local $X$ measurements are used for all three, and in the ideal case a maximally entangled state is produced.
+This is because there exist stabilizer operators consistent with this measurement pattern, as in the exemplary 4 qubit path graph discussed in Section 3.
+Similar structures can be found for all the three graph families mentioned.
+We find sets of stabilizer operators which have $X$ operators in the internal section and Pauli operators which ensure correlations on the terminal qubits – see Section B.2.
+These operators are compatible with the measurement scheme.
+5.2 Bell pair quality scaling for different types of noise
+We analyze the approach of noise-detecting structures for graph families introduced above.
+The procedure is employed for different layer sizes $n$, see Figures 6 and 7.
+We find that the twisted pair and crazy structures are less affected by the noise, as evidenced by the fidelity susceptibility $\alpha$.
+The value of $\alpha$ is smaller, and in the case of the crazy graph does not depend on the graph length.
 Furthermore, the susceptibility for the crazy graph structure does not depend on the length, as a result of the correctional stabilizers mentioned in the previous paragraph.
 This behavior is observed both for uncorrelated phase noise, resulting in probabilistic edge losses and a local $Z$ flip noise.
+The results can be understood as destruction of the perfect correlations present in graph states, mathematically described in Section 3.
+Even if for every ensemble element the end state is maximally entangled, the mixture of such states might have less or no entanglement.
+Postselection on the measurement outcomes consistent with the ideal stabilizer structure helps twofold: one of its effects is probabilistic error detection.
+The other, appearing in the case of the crazy graph ensemble, is more subtle: after postselection, the end state may be the same for a large portion of the ensemble elements (see Appendix B).
+Figure 6: The effects of the edge loss (top) and local phase flip (bottom) noise on the task of Bell pair extraction for the three analyzed graph state families.
+The susceptibility $\alpha$ (Equation 22) is the initial falloff of fidelity in the low noise limit of the resulting two-qubit state with respect to the expected Bell pair.
+The results here only take into account the measurement outcomes consistent with the ideal state.
+In this case (low noise limit) the dominant contributions are the single edge losses (top) and single qubit flips (bottom).
+The constant susceptibility of the crazy graph is a result of the noise at the terminal edges only: each internal edge loss is automatically corrected for and every single flipped qubit is detected.
 An additional entanglement-preserving structure can be found in the case of perfectly correlated phase noise.
-The authors of Ref.
-[13] recently observed that in the case of the path graph, postselection on different measurement outcomes yields vastly different results in the terms of entanglement quality; we have found that this result does generalize to the other linear graphs analyzed by us.
-Both for the crazy and double twist graph families the lowest order nontrivial noise effects can be canceled completely if all the measurement outcomes are postselected on measuring the minus outcome of $X$.
+This is the source of zero susceptibility of twist graph and periodic behavior of crazy graph in Figure 7.
+The authors of Reference [21] recently observed that in the case of the path graph, postselection on different measurement outcomes yields vastly different results in the terms of entanglement quality.
+We have found that this result does generalize to the other linear graphs analyzed by us.
+Both for the crazy and twist graph families the lowest order nontrivial noise effects can be canceled completely if all the measurement outcomes are postselected on measuring the $(-1)$ outcome of $X$.
+Figure 7: The effects of correlated phase noise on the task of Bell pair extraction for the three analyzed graph state families.
+The susceptibility $\alpha$ (Equation 22) is the initial falloff of fidelity in the low noise limit of the resulting two-qubit state with respect to the expected Bell pair.
+Here, only the cases postselected on observing exclusively minus signs during the $X$ measurement are taken into account.
+In this case (low noise limit with restrictive postselection) the twist graph is able to remove the effect of noise completely, and the crazy graph shows a periodic pattern of susceptibilities, still significantly reducing the noise effect.
+The situation inverts for observing exclusively $+$ signs: then, the noise gets amplified.
 This is consistent with our previous observations: such an observed measurement pattern does not violate parity constraints arising from the stabilizer structure.
-This result can be understood by studying the modifications of the stabilizer operators of |$\ket{G}$ under correlated controlled-phase noise.
-For the square graph presented in Figure 1a, if instead of the perfect controlled-$Z$ unitary, the $\operatorname{C\Phi}(\pi+\varepsilon)=\operatorname{diag}(1,1,1,-1\exp(i\varepsilon))$ is used (so that $\operatorname{C\Phi}=\operatorname{CZ}$ for $\varepsilon=0$), the four-qubit entangled state is a weighted graph state |$\ket{G}_{\varepsilon}$.
-If qubits 1 and 3 are measured in the $X$ basis, we expect the measurement outcomes to be equal (we exclude the cases when it does not hold).
-With the weighted graph state |$\ket{G}_{\varepsilon}$ as the initial state, let us denote the unnormalized post-measurement state by
-|$\ket{X_{1,3}\pm}=\frac{\mathbbm{1}\pm X_{1}}{2}\frac{\mathbbm{1}\pm X_{3}}{2}\ket{G}_{\varepsilon}.$ (25)
-The relevant expectation values defining the end correlations (with $\langle O\rangle_{\pm}$ denoting the expectation value of $O$ for the unnormalized state |$\ket{X_{1,3}\pm}$) across qubits $2$ and $4$ are the following, in the lowest nontrivial noise order:
-$\langle Z_{2}Z_{4}\rangle_{+}$ $=1-\varepsilon^{2},$ $\langle Z_{2}Z_{4}\rangle_{-}$ $=-1+\frac{\varepsilon^{2}}{2},$ (26)
-$\langle X_{2}X_{4}\rangle_{+}$ $=1-3\varepsilon^{2},$ $\langle X_{2}X_{4}\rangle_{-}$ $=-1+\frac{\varepsilon^{2}}{2},$
-$\langle Y_{2}Y_{4}\rangle_{+}$ $=-1+3\varepsilon^{2},$ $\langle Y_{2}Y_{4}\rangle_{-}$ $=1+\frac{\varepsilon^{2}}{2},$
-$\||X_{1,3}+\rangle\|^{2}$ $=1-\varepsilon^{2},$ $\||X_{1,3}-\rangle\|^{2}$ $=1-\frac{\varepsilon^{2}}{2}.$
-Note that for the |$\ket{X_{1,3}-}$ outcome the squared length $\||X_{1,3}-\rangle\|^{2}$ is proportional to the unnormalized expectation values: thus, for the physical correlations the dominant noise effects are canceled and only higher order terms ($O(\varepsilon^{4})$) remain.
-On the other hand, if the outcome is |$\ket{X_{1,3}+}$, the noise is amplified compared to postselection on only stabilizer-consistent outcomes.
-Thus, further postselection can amplify the entanglement quality of the remaining qubits at the cost of discarding certain outcomes and reduced production rates.
-The double twist graph can be viewed as multiple such squares stacked together by the terminal vertices: thus, a proper postselection can mitigate the first nontrivial noise effects completely (see Figure 7).
-A similar structure appears for the crazy graph: it is much more complex to analyze it, but computer algebra systems provide evidence of a periodic structure of susceptibilities as the length of the graph increases.
-Figure 6: Susceptibility of the Bell pair entanglement quality to edge loss (top) and local phase flip (bottom) noise.
-Pictured is the linear part of concurrence decrease for the three analyzed families (with postselection on the measurement outcomes consistent with the ideal state) in the low noise limit, where the dominant contributions are the single edge losses (top) and single qubit flips (bottom).
-The constant susceptibility of the crazy graph is a result of the noise at the terminal edges only: each internal edge loss is automatically corrected for and every single flipped qubit is detected.
-Figure 7: Susceptibility of the Bell pair entanglement quality for correlated phase noise.
-Here, only the cases postselected on observing exclusively minus signs during the $X$ measurement are taken into account.
-In this regime, the twist graph is able to remove the effect of noise completely (in the lowest nontrivial orders), and the crazy graph shows a periodic pattern of susceptibilities, still significantly reducing the noise effect.
-The situation inverts for observing exclusively $+$ signs: then, the noise gets amplified.
-6 Extraction of GHZ states
-The methodology applied for Bell pair extraction generalizes to the extraction of GHZ states [23]: generally, for $N$ qubits, they correspond to a star graph with a central qubit connected to the remaining $N-1$ ones (or, equivalently, a complete graph of $N$ qubits).
-Here, we concentrate on the case $N=3$ and adopt star-shaped graph structures built from the graph patterns used in Bell pair creation.
-Refer to Figure 8a for an illustration of the structure: it is built on the template of several crazy graphs, stitched at the central pair of vertices.
-The fidelity response, presented in Figure 8b, is more resilient to noise (as compared to the simpler path graph) also in this new task of extraction of the GHZ state.
-The data presented in this figure require calculations involving a large number of ensemble states and measurement outcomes.
-To mitigate the numerical difficulty, a Monte Carlo method was utilized.
-The figure of merit – fidelity – is linear with respect to the ensemble- and measurement outcome averaging, provided the target state is pure.
-Thus, this approach allows for an approximation of fidelity by sampling ensemble elements and measurement outcomes, avoiding direct calculations with density operators.
-The fidelity of the output state to the ideal pure state is estimated using this method until the plots converge smoothly and the variance of the fidelity estimator is negligible.
-Analysis of the data in Figure 8 shows that the fidelity response for both the GHZ and Bell state extractions exhibits qualitatively similar behavior: most importantly, the susceptibility is greatly reduced in the low noise limit.
-Thus, the ’crazy graph’ structures can generalize beyond Bell pair extraction to more complex multi-qubit systems, maintaining their noise-resilient properties.
-This performance gain is due to the noise-detecting and -correcting capabilities intrinsic to these graph families.
-Further investigations of different connectivity structures suggest that this behavior is universal, although a formal proof is beyond the scope of this study.
-If true, the ’crazy graph’ can serve as a link similar to a simple path link, similar in function but implementing some level of error correction.
+This result can be understood by studying the modifications of the stabilizer operators of |$\ket{G}$ under correlated controlled-phase noise – see Appendix D.
+6 Extraction of GHZ states and large noise regime
 a)Refer to caption
@@
 Figure 8: a) Structures of the analyzed graphs analyzed: path and crazy templates yielding Bell pairs and 3-qubit GHZ states upon measurements of the central qubits.
+Similarly to the previous case, all qubits except the terminal ones (highlighted and enlarged) are measured in the $X$ basis.
 b) Fidelity of the end state with respect to the ideal Bell and GHZ states as a function of noise.
@@
 The initial fall-off rate of fidelity at $p=0$ is the fidelity susceptibility, defined in Equation 22.
+In Section 5 we concentrated on the Bell pair extraction for various geometries in the low noise regimes.
+However, the presented structures (twisted and crazy graphs) are of more general use.
+In this section, we analyze large noise effects on two exemplary scenarios.
+In addition to Bell pair extraction, we consider a similar procedure for GHZ states [35] arising from a starlike graph.
+The considered graph structures are presented in Figure 8a.
+The main result here is the fidelity of output states (Figure 8b) postselected on observing the expected sign structures during $X$ measurement of the internal qubits.
+The Monte Carlo simulations required to analyze the large noise regime are lengthy, as multiple ensemble elements are relevant for the result.
+Therefore, for simplicity we consider fixed graph geometries (path and crazy-like with fixed length) with only the uncorrelated phase noise present.
+The fidelity $\mathcal{F}$ is determined as the average of ensemble elements and postselected stabilizer-consistent outcomes – see Equation 21.
+This avoids direct calculations with density operators.
+The fidelity of the output state to the ideal pure state is estimated using this method until the plots converge smoothly and the variance of the fidelity estimator is negligible.
+Analysis of the Figure 8 shows that the fidelity response for both the GHZ and Bell state extractions exhibits qualitatively similar behavior.
+The output state is more resilient to noise (as compared to the simpler path graph) also in this new task of extraction of the GHZ state.
+Most importantly, the noise effect is greatly reduced in the intermediate noise limit.
+This performance gain is due to the postselection involving noise-detecting and -correcting capabilities intrinsic to this graph family.
+Thus, the ‘crazy graph’ structures can serve as a general link.
+They can generalize beyond Bell pair extraction to more complex multi-qubit systems, maintaining their noise resistant properties.
+Further investigations of different connectivity structures suggest that this behavior is universal, although a formal proof is beyond the scope of this study.
+If true, the ‘crazy graph’ can serve as a link similar to a simple path link.
+That is, it has similar in functions but some level of error correction is implemented.
 7 Summary and Outlook
@@
 This approach uses additional qubits for probabilistic verification and demonstrates a new method to improve resilience against specific noise disturbances.
+The mathematical formalism used in the calculations has been previously used in the nonlocality detection in graph states [52, 53] and could be extended to other families beyond Bell and GHZ states.
 Future research may include more general tasks found in the measurement-based quantum computation approach.
-So far, our results are applicable to measurement patterns composed of Pauli operators, but it is known to restrict the set of possible operations to a clasically simulable one: a similar method, based on non-Pauli measurements, but still allowing for noise effects reduction, could open new pathways for quantum computation, albeit with conceptual and practical complexities.
+So far, our results are applicable to measurement patterns composed of Pauli operators, but it is known to restrict the set of possible operations to a classically simulable one.
+A similar method, based on non-Pauli measurements, but still allowing for noise effects reduction, could open new pathways for quantum computation, albeit with conceptual and practical complexities.
 The methods presented here can be generalized for arbitrary output graph geometry, and the applicability of this procedure for quantum metrology [54] could be studied.
@@
 16KIS1618K), the Stiftung der Deutschen Wirtschaft, and the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement (No 945422).
-Appendix A Post-measurement correlations and stabilizers compatible with the measurement scheme
+Funding from the projects DeQHOST APVV-22-0570, QUAS VEGA 2/0164/25, and Stefan Schwarz programme is gratefully acknowledged.
+Appendix A Bell pair creation is equivalent to teleportation
+As mentioned in the introduction, the creation of a Bell pair is equivalent to the task of information transfer (teleportation) across the graph.
+This is so because if the Bell pair creation in the graph $G$ is viewed in the context of a larger graph $G^{\prime}$, the entire process, followed by measurement of one of the terminal qubits, performs qubit fusion: the unmeasured terminal qubit acts as the two combined.
+Since all local measurements with nonoverlapping supports do commute, this can be done before any other measurement or afterwards: in the latter case, the information created in one of the qubits is merged with the second one.
+This is captured by the following Lemma, represented pictorially by the Figure 9.
+Figure 9: If a measurement scheme generates a Bell pair across the vertices $\{{\operatorname{in}},{\operatorname{out}}\}$ in the state corresponding to a graph $G$, in a larger $G^{\prime}$ it can be used to transfer information across the graph.
+Extending the measurement by a properly chosen operator $A_{{\operatorname{in}}}$ leads to the fusion of the two terminal vertices.
 Lemma 3.
-Given a graph state |$\ket{G}$ of $N$ qubits and a certain measurement scheme defined in Lemma 1 by a set of qubits $I$ and operators $o_{i}$, then the subset of the stabilizer of |$\ket{G}$ compatible with the measurement pattern has $N-\lvert I\rvert$ generators, and therefore fully defines the post-measurement state.
+Let us assume the following:
+  • We fix a graph $G=\{V,E\}$, and choose the two terminal vertices $\{{\operatorname{in}},{\operatorname{out}}\}$ such that $V=\{{\operatorname{in}},{\operatorname{out}}\}\cup I$.
+  • The corresponding graph state |$\ket{G}$ undergoes a sequential measurement of $I$ producing a Bell pair across $\{{\operatorname{in}},{\operatorname{out}}\}$.
+According to the Lemma 1, this means that the correlations between $\{{\operatorname{in}},{\operatorname{out}}\}$ are described by stabilizer operators of |$\ket{G}$ modified by the measurement process (Equation 12).
+  • Let $G$ be a part of a larger graph $G^{\prime}$ in the following sense: $G^{\prime}=(V^{\prime},E^{\prime})$ embeds $G$ in such a way that $\{{\operatorname{in}},{\operatorname{out}}\}$ is a separating set.
+This means that the only connections between $G$ and the rest of $G^{\prime}$ are at these two vertices.
+  • Let us denote the vertices in $V^{\prime}\setminus V$ to which both qubits are connected by $\mathcal{N}^{\prime}({\operatorname{in}})$ and $\mathcal{N}^{\prime}({\operatorname{out}})$, respectively, and assume that the two sets are disjoint.
+Consider now the following process viewed from the perspective of the graph $G^{\prime}$:
+  1. (a)
+    the measurement scheme, which within $G$ would produce a Bell pair across the terminal qubits, followed by
+  2. (b)
+    a product of single qubit unitary operations on the qubits $\{{\operatorname{in}},{\operatorname{out}}\}$, aimed to counteract the effect of measurement and bring back the stabilizer operators into a manageable form, and
+  3. (c)
+    the measurement of $X_{\operatorname{in}}$.
+These operations taken together effectively performs qubit fusion of the terminal qubits.
+The new state is a graph state where the qubits $V^{\prime}\setminus V$ are intact, and the new neighborhood of the remaining qubit is $\mathcal{N}_{{\operatorname{out}}}\cup\mathcal{N}_{{\operatorname{in}}}$.
+Proof.
+If within $G$ a Bell pair is produced across the terminal vertices $\{{\operatorname{in}},{\operatorname{out}}\}$ as a result of the measurement pattern $O=\prod_{i\in I}o_{i}$, this means that there exist three $O$-consistent stabilizers involving the terminal vertices in the sense of Lemma 1.
+Combinatorical considerations show that one of those must be of a form such that the operators at the terminal vertices are $A_{\operatorname{in}},B_{\operatorname{out}}\in\{X,Y\}$ (but not $Z$), chosen independently.
+Thus, there exists an $G$-stabilizer operator of the form ±$\pm A_{\operatorname{in}}B_{\operatorname{out}}\prod_{j\in J}o_{j}=g_{\operatorname{in}}g_{\operatorname{out}}\prod_{j\in J^{*}}g_{j}$, with $J,J^{*}\subset I$.
+Note that the definition of $g_{i}$ depends on the graph in question (Equation 6): if by $g^{\prime}_{i}$ we denote the generators corresponding to the entire graph $G^{\prime}$, the result is also a stabilizer operator, differing only by the $Z$ operators in the set $\mathcal{N}^{\prime}({\operatorname{in}})\cup\mathcal{N}^{\prime}({\operatorname{out}})$.
+Let us now determine the post-measurement stabilizer operator within $G^{\prime}$ by application of Lemma 1.
+As prescribed by the operator $O$, every qubit $i\in L$ is measured in the basis of $o_{i}$, and the qubit $({\operatorname{in}})$ is measured in the basis of $A_{\operatorname{in}}$.
+As a result, the post-measurement stabilizer operator is
+(±$\left(\pm s_{\operatorname{in}}\prod_{j\in J}s_{j}\right)B_{\operatorname{out}}\mathchoice{}{}{}{}\prod_{j^{\prime}\in\mathcal{N}^{\prime}_{\operatorname{in}}\cup\mathcal{N}^{\prime}_{\operatorname{out}}}\mathchoice{}{}{}{}Z_{j^{\prime}},$ (23)
+which can be taken to be a new generator associated with the vertex ${\operatorname{out}}$, and brought to the canonical form of $X_{\operatorname{out}}\prod_{j^{\prime}\in\mathcal{N}^{\prime}_{\operatorname{in}}\cup\mathcal{N}^{\prime}_{\operatorname{out}}}Z_{j^{\prime}}$ by local unitary basis change of the ${\operatorname{out}}$-qubit.
+The properties of other stabilizers involving the qubits $\{{\operatorname{in}},{\operatorname{out}}\}$ can be proven similarly.
+They merge to stabilizer operators of only the qubit remaining out-qubit.
+Thus, after the measurement within the Bell-generating part $I\cup\{{\operatorname{in}}\}$, the ${\operatorname{out}}$-qubit behaves exactly like it was connected to the neighborhood $\mathcal{N}^{\prime}_{\operatorname{in}}$ as well: the two vertices are fused and the information between them is transferred.
+∎
+Appendix B Post-measurement correlations and stabilizers compatible with the measurement scheme
+Lemma 4.
+Fix a graph state |$\ket{G}$ of $N$ qubits and a certain measurement scheme defined in Lemma 1 by a set of qubits $I$ and operators $o_{i}$.
+The stabilizer operators of |$\ket{G}$ which are compatible with the measurement pattern (that is, the operators share elements with the measurement scheme, see discussion before Equation 12) form a subgroup.
+This subgroup has $N-\lvert I\rvert$ generators, and therefore fully defines the post-measurement state.
 Proof.
@@
 The identification is following: if the stabilizer operator is defined as $\prod_{j\in J}g_{j}$, the corresponding element of $\mathbb{Z}_{2}^{2N}$ is $(x_{1},\ldots,x_{N})\oplus(z_{1},\ldots,z_{N})$, with $x_{i}$ being 1 if and only if $i\in J$ and $\vec{z}=A_{G}\vec{x}\mod 2$, where $A_{G}$ is the adjacency matrix of $G$.
-Thus, if a stabilizer contains $X_{i}$, the corresponding numbers are $(x_{i},z_{i})=(1,0)$, for $Z_{i}$ it is $(x_{i},z_{i})=(0,1)$, and $Y_{i}$ corresponds to $(x_{i},z_{i})=(1,1)$; no operator at site $i$ is denoted by $(x_{i},z_{i})=(0,0)$.
-In this way, the additive algebra of $\mathbb{Z}_{2}^{2}$ mimics the product rules of Pauli operators with only the missing sign structure.
+Thus, if a stabilizer operator contains $X_{i}$, the corresponding numbers are $(x_{i},z_{i})=(1,0)$, for $Z_{i}$ it is $(x_{i},z_{i})=(0,1)$, and $Y_{i}$ corresponds to $(x_{i},z_{i})=(1,1)$; no operator at site $i$ is denoted by $(x_{i},z_{i})=(0,0)$.
+In this way, the additive algebra of $\mathbb{Z}_{2}^{2}$ mimics the product rules of Pauli operators with only the sign structure missing.
 The stabilizer of |$\ket{G}$ corresponds to a special $N$-dimensional subspace in $\mathbb{Z}_{2}^{2N}$ [7].
-For any stabilizer operator to be compatible with the measurement scheme in the sense of Lemma 1, each operator of its Pauli string form with support in $I$ must be $o_{i}$: this introduces constraints to the set of compatible operators.
-However, these constraints translated into the language of $\mathbb{Z}_{2}^{2N}$ are linear: if $X_{i}$ is measured at site $i$, the additional constraint is $z_{i}=0$ (this encompasses the cases of $X_{i}$ and identity at site $i$), if $Z_{i}$ is measured, the constraint is $x_{i}=0$, and for $Y_{i}$ it is $x_{i}+z_{i}=0\mod 2$.
+For any stabilizer operator $Q=\prod_{j\in J}q_{j}$ to be compatible with the measurement scheme $O=\prod_{i\in I}o_{i}$ in the sense of Lemma 1, each operator of $q_{i}$ its Pauli string form with support in $I$ must be $o_{i}$: this introduces constraints to the set of compatible operators.
+These constraints translated into the language of $\mathbb{Z}_{2}^{2N}$ are linear: if $X_{i}$ is measured at site $i$, the additional constraint is $z_{i}=0$ (this encompasses the cases of $X_{i}$ and identity at site $i$), if $Z_{i}$ is measured, the constraint is $x_{i}=0$, and for $Y_{i}$ it is $x_{i}+z_{i}=0\mod 2$.
 All of them are independent, and there are $\lvert I\rvert$ of them, one for each measured site.
@@
 For example, the correlations that stem from the measurement described after Lemma 1 and in Figure 3 correspond to the following $\vec{x}\oplus\vec{z}$ vectors in the binary representation:
-$(0,1,0,1,0)$ $\oplus(1,0,0,0,1),$ (27)
-$(1,0,1,0,1)$ $\oplus(0,0,0,0,0),$
-$(1,1,1,1,1)$ $\oplus(1,0,0,0,1).$
+[…]
+[…]
+[…]
 The third vector is a linear combination of the first two; thus the space of end correlations is two-dimensional.
-A.1 Patterns for specific graph families
+B.1 Discussion of measurement outcomes
+Figure 10: Some measurement patterns have a restricted set of accepted outcomes.
+If we measure the pairs of qubits associated to the vertices in the gray areas in Pauli $X$ basis, the products of the outcomes have to be equal to one.
+This indicates which combination of signs are allowed: The accepted signs are shown in the green areas while the rejected signs are shown in the red areas.
+We postselect by discarding all states whose measurement outcomes have signs which disagree with the target graph state.
+[…]
+In Sections 5 and 6, we perform measurements on noisy graph states and postselect on measurement outcomes.
+That is, we consider all stabilizer operators with support within the measurement pattern and check the outcome for consistency.
+If the outcomes do not match the expectation, we discard the resulting state.
+This concept is also pictured in Figure 2.
+B.2 Patterns for specific graph families
 Consider a path graph $G$ of $n+2$ vertices $V=\{0,1,\ldots,n,n+1\}$ with edges $E=\{\{k,k+1\}\}_{k=0}^{n}$.
@@
 $Z_{0}X_{n+1}\prod_{k\in I}s_{k},$ $\leavevmode\nobreak\ X_{0}Z_{n+1}\prod_{k\in I^{\prime}}s_{k}.$ (even $(\text{even }n)$
-Both cases are consistent with a maximally entangled two-qubit state being produced in the terminal qubits: for the odd-$n$ case, the existence of stabilizers described by Equation 30 ensures that further measurement of $X$, $Y$ or $Z$ at qubit 00 is perfectly (anti)correlated with the outcome of measurement of the same operator at qubit $n+1$.
+Both cases are consistent with a maximally entangled two-qubit state being produced in the terminal qubits: for the odd-$n$ case, the existence of stabilizers described by LABEL:eqn:termoddpathcorr ensures that further measurement of $X$, $Y$ or $Z$ at qubit 00 is perfectly (anti)correlated with the outcome of measurement of the same operator at qubit $n+1$.
 This result generalizes to the case where each qubit $k$ is replaced by a set of $m_{k}$ qubits in such a way that the adjacent layers are fully connected.
@@
 This class of graphs includes the path graph (each layer has only a single qubit), the crazy graph (the layers 00 and $n+1$ contain one qubit each, every other has two) and twisted graphs (the qubit count alternates between 1 and 2) as well.
-Suppose that all the internal qubits in every layer are measured in the $X$ basis.
+Suppose that all the qubits in every internal ($i=1,\ldots,n$) layer are measured in the $X$ basis.
 The stabilizer operators that determine the remaining correlations must be consistent with this structure.
-A particularly simple set of such operators exists: first, observe that if a layer $k$ contains at least two qubits, $X_{k,i}X_{k,i^{\prime}}$ is a stabilizer operator.
+A particularly simple set of such operators exists: first, observe that if a layer $k$ contains at least two qubits $i$ and $i^{\prime}$, then $X_{k,i}X_{k,i^{\prime}}$ is a stabilizer operator.
 Thus, the signs of outcomes have to be equal within a layer.
-Therefore, for any choice of inner-layer qubit index $(i_{k})$ the direct analogue of Equation 28 holds:
+Therefore, for any choice of inner-layer qubit index $(i_{k})$ the direct analogue of LABEL:eqn:stabextended holds:
 $Z_{0}Z_{n+1}\prod_{k\in I}X_{k,i_{k}},$ $\leavevmode\nobreak\ X_{0}X_{n+1}\prod_{k\in I^{\prime}}X_{k,i_{k}},$ (odd $(\text{odd }n),$ (29)
@@
 Thus, the terminal layers 00 and $n+1$ always consist only of one qubit each, and the inner indices are omitted.
-Consequently, if $s_{k}$ in Equation 30 are interpreted as signs of outcomes at layer $k$ (which are all equal within a layer due to the structure of the stabilizer operators mentioned previously), the same post-measurement stabilizer structure appears.
-In the case of the crazy graph, an additional observation can be made: loss of any single internal (not adjacent to the terminal qubits) edge yields the same post-measurement stabilizer structure, since the indices $(i_{k})$ in Equation 32 can be chosen arbitrarily.
+Consequently, if $s_{k}$ in LABEL:eqn:termoddpathcorr are interpreted as signs of outcomes at layer $k$ (which are all equal within a layer due to the structure of the stabilizer operators mentioned previously), the same post-measurement stabilizer structure appears.
+In the case of the crazy graph, an additional observation can be made: loss of any single internal (not adjacent to the terminal qubits) edge yields the same post-measurement stabilizer structure, since the indices $(i_{k})$ in LABEL:eqn:stabextendedgeneral can be chosen arbitrarily.
 Thus, the crazy graph family is especially tolerant to single edge loss: it does not affect the post-measurement state, if the outcomes are postselected on the ideal sign structure.
 This can be generalized to the loss of any subset of edges not adjacent to each other and the terminal qubits.
-Appendix B Derivation for the noise models
-B.1 Uncorrelated phase noise
+Appendix C Derivation for the noise models
+C.1 Uncorrelated phase noise
 Figure 11: The controlled-$Z$ gate, if implemented as controlled phase operation, is susceptible to phase variations.
@@
 For just a single edge, if the phase distribution is centered around $\pi$ (corresponding to the ideal graph state), the resulting quantum channel is a mixture of controlled-$Z$ and identity: the edge is probabilistically generated.
-In certain ion traps and superconducting quantum devices, the sequence of $\operatorname{CZ}$ gates is realized with evolution with an Ising-type Hamiltonian:
+In certain ion traps and superconducting quantum devices, the sequence of $\operatorname{CZ}$ gates is realized with evolution with an Ising-type Hamiltonian [41, 42, 43, 44] :
 $H=\sum_{(i,j)\in G}\gamma_{i,j}\frac{Z_{i}+1}{2}\frac{Z_{j}+1}{2}.$ (30)
@@
 $p=\frac{1-\exp\left(-\frac{\sigma^{2}}{2}\right)}{2}=\frac{\sigma^{2}}{4}-\frac{\sigma^{4}}{16}+O(\sigma^{6}).$ (34)
-B.2 Correlated phase noise
+C.2 Correlated phase noise
 Correlated phase noise has similar, but more involved effect.
@@
 Thus, in the lowest order or approximation the output state is described by:
-  • The unmodified state graph state |$\ket{G}$ with reduced probability coming from $\operatorname{\mathcal{CZ}}$ terms appearing in product of channels defined by Equation 34,
-  • graph states with one edge missing as in the case of uncorrelated noise,
+  • The unmodified state graph state |$\ket{G}$, with reduced probability coming from $\operatorname{\mathcal{CZ}}$ terms appearing in product of channels defined by Equation 31,
+  • graph states with one edge missing, as in the case of uncorrelated noise,
   • graph states modified by a product of two $\operatorname{\mathcal{CS}}^{\pm}$ channels corresponding to different edges, with possible negative weights coming from negative signs in Equation 31.
@@
 $\rho=$ $|\displaystyle(1-p\lvert E\rvert)\ket{G}\bra{G}+p\sum_{e\in E}\lvert{\!\!\overbrace{\!G^{\prime}\!}^{G\setminus\{e\}}\!\!}\rangle\bra{G^{\prime}}$ | (35)
-$+\frac{p}{2}\sum_{(e,e^{\prime}\in E}^{e\neq e^{\prime}}\sum_{\pm_{1},\pm_{2}}\pm_{1}\pm_{2}\operatorname{\mathcal{CS}}_{e}^{\pm_{1}}\operatorname{\mathcal{CS}}_{e^{\prime}}^{\pm_{2}}[\ket{G}\bra{G}].$
+$+\frac{p}{2}\sum_{(e,e^{\prime}\in E}^{e\neq e^{\prime}}\sum_{s_{1},s_{2}=\pm 1}s_{1}s_{2}\operatorname{\mathcal{CS}}_{e}^{s_{1}}\operatorname{\mathcal{CS}}_{e^{\prime}}^{s_{2}}[\ket{G}\bra{G}].$
 Here, $\lvert E\rvert$ denotes the number of edges in the graph $G$.
-Note that it is not an ensemble decomposition: in the last sum, some signs are negative.
+Note that this decomposition is not convex and cannot be interpreted as an ensemble: the minus signs in front of the decomposition terms prevent this.
+Signs appear as a result of discrete decomposition of controlled-phase and subsequent averaging; see Appendix C for details.
+In this decomposition, $G-e\coloneqq(V,E\setminus\{e\})$ is the graph with edge $e$ removed, and ${\operatorname{\mathcal{CS}}^{\pm}}$ is the application of one of the two unitary square roots of $\operatorname{CZ}$:
+${\operatorname{\mathcal{CS}}^{\pm}}[\rho]=\sqrt{\operatorname{CZ}^{\pm}}\rho\sqrt{\operatorname{CZ}^{\pm}}^{\dagger},$ (36)
+where
+$\sqrt{\operatorname{CZ}^{\pm}}\coloneqq\frac{1}{2}\left[\left(1+e^{\pm i\frac{\pi}{2}}\right)\operatorname{\mathds{1}}+\left(1-e^{\pm i\frac{\pi}{2}}\right)\operatorname{CZ}\right]=\begin{pmatrix}1&&&\\ &1&&\\ &&1&\\ &&&\pm i\end{pmatrix}.$ (37)
 This decomposition is not unique, and in principle different weights for the full, edge-missing, and $\operatorname{\mathcal{CS}}$-modified graph states are possible; however, some of the signs will always be negative, regardless of the choice.
-However, such decompositions are only low noise approximations for $p\ll\lvert E\rvert^{-1}$.
-B.3 Fusion gate noise
+All such decompositions are only low noise approximations for $p\ll\lvert E\rvert^{-1}$.
+C.3 Fusion gate noise
 Figure 12: Polarizing beamsplitter transmits horizontally polarized photons and reflects vertically polarized ones, entangling the input state in the output modes.
 Subsequent polarization measurement in the diagonal basis of one of the output modes is a basis for type 1 fusion gate, which is used in generation of larger graph states from Bell pairs.
-Optical realizations [41], while using different mechanisms for entangling, lead to a similar result.
-If a type 1 entangling gate is used, imperfect mode matching leads affected qubits to be effectively depolarized in the $Z$ basis.
-This has the exact same result on the output state as local (single-qubit) phase noise, despite different physical origins.
-A different scheme for generating large cluster states is employed in linear optics experiments.
+A different scheme for generating large cluster states is employed in linear optics experiments [56].
 In such systems, the qubits are encoded as photons in different light modes.
 Two modes are associated with each qubit, and a single photon being in one of them corresponds to the standard computational basis states.
-Usually, these are polarization (horizontal/vertical) modes with overlapping spatial profiles.
+Then, the graph states are built by qubit fusion of initially generated Bell pairs.
+If a type 1 entangling gate is used for this, imperfect mode matching leads affected qubits to be effectively depolarized in the $Z$ basis.
+This has the exact same result on the output state as local (single-qubit) phase noise, despite different physical origins.
+Usually, the two modes defining qubit degrees of freedom are polarization (horizontal/vertical) modes with overlapping spatial profiles.
 Thus, let us denote the creation operators of those modes by ${\hat{h}^{\dagger}}_{i}$ and ${\hat{v}^{\dagger}}_{i}$.
@@
 |$\ket{G}$ =$=\prod_{\{1,j\}\in E}\operatorname{CZ}_{1,i}\prod_{\{2,j\}\in E}\operatorname{CZ}_{2,j}\ket{+}\otimes\ket{+}\otimes\ket{G^{\prime}}$ (40)
-$=\frac{{\hat{h}^{\dagger}}_{1}{\hat{h}^{\dagger}}_{2}f_{++}+{\hat{v}^{\dagger}}_{1}{\hat{h}^{\dagger}}_{2}f_{-+}+{\hat{h}^{\dagger}}_{1}{\hat{v}^{\dagger}}_{2}f_{+-}+{\hat{v}^{\dagger}}_{1}{\hat{v}^{\dagger}}_{2}f_{--}}{2}\ket{\emptyset},$
-where the symbols $f$ denote expressions involving creation operators to arrive at the state |$\ket{\psi}$.
-The $f_{++}$ prepares the graph state |$\ket{G^{\prime}}$, while $f_{-+}$ creates the state $\prod_{(1,i)\in G}z_{i}\ket{G^{\prime}}$: it is the postselected graph state upon measuring negative sign of the first qubit in the $Z$ basis.
-The other two operators work similarly; however, $f_{--}$ flips the phase of |$\ket{G^{\prime}}$ only in the qubits that are connected to only $1$ or $2$.
-Polarizing beamsplitter exchanges the vertical polarization, keeping the horizontal ones intact ($U{\hat{h}^{\dagger}}_{1,2}={\hat{h}^{\dagger}}_{1,2}U$, $U{\hat{v}^{\dagger}}_{1,2}={\hat{v}^{\dagger}}_{2,1}U$):
-$U\ket{G}$ $=\frac{{\hat{h}^{\dagger}}_{1}{\hat{h}^{\dagger}}_{2}f_{++}+{\hat{v}^{\dagger}}_{2}{\hat{h}^{\dagger}}_{2}f_{-+}+{\hat{h}^{\dagger}}_{1}{\hat{v}^{\dagger}}_{1}f_{+-}+{\hat{v}^{\dagger}}_{1}{\hat{v}^{\dagger}}_{2}f_{--}}{2}\ket{\emptyset}.$ (42)
+$=\frac{{\hat{h}^{\dagger}}_{1}{\hat{h}^{\dagger}}_{2}\hat{f}_{++}+{\hat{v}^{\dagger}}_{1}{\hat{h}^{\dagger}}_{2}\hat{f}_{-+}+{\hat{h}^{\dagger}}_{1}{\hat{v}^{\dagger}}_{2}\hat{f}_{+-}+{\hat{v}^{\dagger}}_{1}{\hat{v}^{\dagger}}_{2}\hat{f}_{--}}{2}\ket{\emptyset},$
+where the symbols $\hat{f}$ denote expressions involving creation operators to arrive at the state |$\ket{\psi}$.
+The $\hat{f}_{++}$ prepares the graph state |$\ket{G^{\prime}}$, while $\hat{f}_{-+}$ creates the state $\prod_{(1,i)\in G}z_{i}\ket{G^{\prime}}$: it is the postselected graph state upon measuring negative sign of the first qubit in the $Z$ basis.
+The other two operators work similarly; however, $\hat{f}_{--}$ flips the phase of |$\ket{G^{\prime}}$ only in the qubits that are connected to only $1$ or $2$.
+Polarizing beamsplitter, described mathematically by a unitary operator $U$, exchanges the vertical polarization, keeping the horizontal ones intact ($U{\hat{h}^{\dagger}}_{i}={\hat{h}^{\dagger}}_{i}U$, $U{\hat{v}^{\dagger}}_{1}={\hat{v}^{\dagger}}_{2}U$, and $U{\hat{v}^{\dagger}}_{2}={\hat{v}^{\dagger}}_{1}U$).
+First part of the fusion process leads to:
+$U\ket{G}$ $=\frac{{\hat{h}^{\dagger}}_{1}{\hat{h}^{\dagger}}_{2}\hat{f}_{++}+{\hat{v}^{\dagger}}_{2}{\hat{h}^{\dagger}}_{2}\hat{f}_{-+}+{\hat{h}^{\dagger}}_{1}{\hat{v}^{\dagger}}_{1}\hat{f}_{+-}+{\hat{v}^{\dagger}}_{1}{\hat{v}^{\dagger}}_{2}\hat{f}_{--}}{2}\ket{\emptyset}.$ (41)
 The spatial mode 2 is subsequently measured in the basis of $({\hat{h}^{\dagger}}_{2}\pm{\hat{v}^{\dagger}}_{2})\ket{\emptyset}$, corresponding to the $\ket{\pm}$ of the logical qubit.
 If only a single $h+v$ photon is detected, the postselected output state reads
-$\frac{{\hat{h}^{\dagger}}_{1}f_{++}+{\hat{v}^{\dagger}}_{1}f_{--}}{\sqrt{2}}\ket{\emptyset}.$ (43)
+$\frac{{\hat{h}^{\dagger}}_{1}\hat{f}_{++}+{\hat{v}^{\dagger}}_{1}\hat{f}_{--}}{\sqrt{2}}\ket{\emptyset}.$ (42)
 This is exactly the state corresponding to a graph $G^{\prime\prime}=(V^{\prime\prime},E^{\prime\prime})$ for which vertices 1 and 2 were merged with duplicate edges removed: $E^{\prime\prime}=E^{\prime}\cup\{\{1,i\}:\{1,i\}\in E\veebar\{2,i\}\in E\}$.
@@
 Shall this assumption not be met, the interference required for the fusion to work does not take place, and effectively the two photons originating in mode 1 and 2 are measured independently.
-If only one photon was observed in the new mode 2, half of the times it arrived from the mode (1,v), while the photon from mode 2 is now localized in the new mode (1,v) and the final state is ${\hat{v}^{\dagger}}_{1}f_{--}\ket{\emptyset}$.
-The opposite process means that the resulting state ${\hat{h}^{\dagger}}_{1}f_{++}\ket{\emptyset}$.
+If only one photon was observed in the new mode 2, half of the times it arrived from the mode (1,v), while the photon from mode 2 is now localized in the new mode (1,v) and the final state is ${\hat{v}^{\dagger}}_{1}\hat{f}_{--}\ket{\emptyset}$.
+The opposite process means that the resulting state ${\hat{h}^{\dagger}}_{1}\hat{f}_{++}\ket{\emptyset}$.
 This effectively creates an ensemble which allows for the following interpretation: the fusion took place as if the photons were indistinguishable, but it was immediately measured in the $Z_{1}$ basis and the result of measurement is not known.
-This is equivalent to a simple probabilistic $Z_{1}$ flip:
-$\overbrace{\Pi_{Z_{1}+}}^{\frac{1}{2}(1+Z_{1})}\rho\Pi_{Z+}+\overbrace{\Pi_{Z_{1}-}}^{\frac{1}{2}({1-Z_{1}})}\rho\Pi_{Z_{1}-}=\frac{\rho+Z_{1}\rho Z_{1}}{2}.$ (44)
+This is equivalent to a simple probabilistic $Z_{1}$ flip, described by the following channel:
+$\rho\mapsto\overbrace{\Pi_{Z_{1}+}}^{\frac{1}{2}(1+Z_{1})}\rho\Pi_{Z+}+\overbrace{\Pi_{Z_{1}-}}^{\frac{1}{2}({1-Z_{1}})}\rho\Pi_{Z_{1}-}=\frac{\rho+Z_{1}\rho Z_{1}}{2}.$ (43)
 The $Z$-flip ensemble is equivalent to $Z$-measurement one and is easier to work with.
 It is easier to observe that the noise does not propagate if the affected qubit is then fused with others, since there is always just a (potentially $Z$-flipped) graph state at the input – unless it also fails because the noise is correlated.
+Appendix D Cancellation of correlated edge noise effects in twist and crazy graph families
+If for the square graph presented in Figure 2a, instead of the perfect controlled-$Z$ unitary, the controlled phase
+$\operatorname{C\Phi}(\pi+\varepsilon)=\operatorname{diag}(1,1,1,-1\exp(i\varepsilon))$ (44)
+is used with the same $\varepsilon$ for all edges, the result is a four-qubit weighted graph state which we denote by |$\ket{G}_{\varepsilon}$.
+If qubits 1 and 3 are measured in the $X$ basis, we postselect on the measurement outcomes which are equal.
+With the weighted graph state |$\ket{G}_{\varepsilon}$ as the initial state, let us denote the unnormalized post-measurement state by
+|$\ket{\xi}_{\pm}=\frac{\mathbbm{1}\pm X_{1}}{2}\frac{\mathbbm{1}\pm X_{3}}{2}\ket{G}_{\varepsilon}.$ (45)
+Note the equality of signs $\pm$ appearing in two projection operators: we postselect on observing this type of outcome, since unequal signs can not appear in the ideal ($\varepsilon=0$) case.
+The relevant expectation values defining the end correlations (with $\langle O\rangle_{\pm}$ denoting the expectation value of $O$ for the unnormalized state |$\ket{\xi}_{\pm}$) across qubits $2$ and $4$ are the following, in the lowest nontrivial noise order:
+$\langle Z_{2}Z_{4}\rangle_{+}$ $=1-\varepsilon^{2}+O(\varepsilon^{4}),$ $\langle Z_{2}Z_{4}\rangle_{-}$ $=-1+\frac{\varepsilon^{2}}{2}+O(\varepsilon^{4}),$
+$\langle X_{2}X_{4}\rangle_{+}$ $=1-3\varepsilon^{2}+O(\varepsilon^{4}),$ $\langle X_{2}X_{4}\rangle_{-}$ $=-1+\frac{\varepsilon^{2}}{2}+O(\varepsilon^{4}),$
+$\langle Y_{2}Y_{4}\rangle_{+}$ $=-1+3\varepsilon^{2}+O(\varepsilon^{4}),$ $\langle Y_{2}Y_{4}\rangle_{-}$ $=1-\frac{\varepsilon^{2}}{2}+O(\varepsilon^{4}),$
+$\||\xi\rangle_{+}\|^{2}$ $=1-\varepsilon^{2}+O(\varepsilon^{4}),$ $\||\xi\rangle_{-}\|^{2}$ $=1-\frac{\varepsilon^{2}}{2}+O(\varepsilon^{4}).$ (46)
+Note that for the |$\ket{\xi}_{-}$ outcome the squared length $\||\xi\rangle_{-}\|^{2}$ is proportional to the unnormalized expectation values: thus, for the physical correlations the dominant noise effects are canceled and only higher order terms ($O(\varepsilon^{4})$) remain.
+On the other hand, if the outcome is |$\ket{\xi}_{+}$, the noise is amplified compared to postselection on only stabilizer-consistent outcomes.
+Thus, further postselection can amplify the entanglement quality of the remaining qubits at the cost of discarding certain outcomes and reduced production rates.
+The double twist graph can be viewed as multiple such squares stacked together by the terminal vertices: thus, a proper postselection can mitigate the first nontrivial noise effects completely (see Figure 7).
+A similar structure appears for the crazy graph: it is much more complex to analyze it, but computer algebra systems provide evidence of a periodic structure (also visible Figure 7) in of susceptibilities as the length of the graph increases.